### 04 Mar 2026 v1

# A Non-Abelian Approach to Riemann Surfaces

Abstract.
We develop a non-abelian, gauge-theoretic framework for the Schwarzian derivative and for second-order differential equations on Riemann surfaces. As applications, we extend Dedekind’s Schwarzian approach to elliptic periods to generic one-parameter families of curves of genus $g$ by replacing the non-canonical scalar Picard–Fuchs equation of order $2g$ with a canonical second-order equation with $g\times g$ matrix coefficients on the Hodge bundle. In higher dimensions, we discuss periods of a one-parameter family of cubic threefolds via the intermediate Jacobian. Finally, we discuss mass–spring systems in mechanics as a natural testing ground for the non-abelian Schwarzian viewpoint.
footnotetext: I would like to thank Amir Jafari.
“Everyone knows what a curve is until he has studied enough mathematics.”
— Felix Klein
1. Introduction
In a letter to Borchardt published in 1877, Dedekind discovered a striking relation between the Schwarzian derivative and ratios of elliptic integrals [5, §§ 5–6]. One considers the periods of an elliptic curve as functions of the normalized Hauptmodul $j$ on the modular curve $\mathbb{H}/\operatorname{SL}_{2}(\mathbb{Z})$. These periods satisfy a second-order Picard–Fuchs equation (see, e.g., [15, 14])
$\frac{d^{2}y}{dj^{2}}+\frac{1}{j}\frac{dy}{dj}+\frac{31j-4}{144j^{2}(1-j)^{2}}y=0.$
If $\omega_{1}(j),\omega_{2}(j)$ are two independent solutions, their ratio
$\tau(j)=\frac{\omega_{2}(j)}{\omega_{1}(j)}$
is a local uniformizing parameter on $\mathbb{H}$. The ordered pair $(\omega_{1},\omega_{2})$ is determined only up to the natural action of $\operatorname{GL}_{2}(\mathbb{C})$ on the two-dimensional space of solutions; hence $\tau$ is determined only up to a Möbius transformation. Recall that for a locally univalent meromorphic function $f$ one defines the Schwarzian derivative by
$S(f):=\left(\frac{f^{\prime\prime}}{f^{\prime}}\right)^{\prime}-\frac{1}{2}\left(\frac{f^{\prime\prime}}{f^{\prime}}\right)^{2}.$
Following Klein’s account in [16, Chapter III, §6], this invariant goes back to Lagrange (see [17]). Since the Schwarzian derivative is invariant under Möbius transformations, the quantity $S(\tau)$ is well defined, independent of the chosen basis. Dedekind’s computation asserts that $S(\tau)$ is in fact the rational function
$S(\tau)(j)=\frac{3}{8(1-j)^{2}}+\frac{4}{9j^{2}}+\frac{23}{72j(1-j)}.$
From Schwarzian to curvature. A guiding principle of this paper is that Dedekind’s identity is best understood as a curvature statement for a connection naturally attached to the differential equation. Concretely, the pre-Schwarzian $A(f)=f^{\prime\prime}/f^{\prime}$ behaves like a Maurer–Cartan (quantum) connection, and its curvature $F_{A}=A^{\prime}-A^{2}$ is the Schwarzian derivative. On a general Riemann surface the curvature fails to be a genuine quadratic differential: it transforms with a controlled projective anomaly given by the Schwarzian of the coordinate changes. This anomaly is precisely the transformation law of projective connections [11].
The main lemma and theorem show that if $g$ is an invertible matrix-valued solution of $g^{\prime\prime}=2Ag^{\prime}+qg$, then under the gauge action, the connection $g^{-1}\bullet A$ has curvature
$F_{g^{-1}\bullet A}=g^{-1}(F_{A}-q)g.$
As a corollary, the conjugacy invariants of $F_{A}-q$ (traces of powers, coefficients of characteristic polynomials) define invariants of the differential equation; we call them characteristic invariants by analogy with Chern’s characteristic forms [2]. In the scalar case this reduces to Dedekind’s classical mechanism: the Schwarzian of the ratio of periods is determined by the coefficients of the Picard–Fuchs equation.
Applications:
We discuss the following three applications of the non-abelian Schwarzian formalism.
  1. (i)
    Periods of Riemann surfaces of higher genus (Sections 12 and 13). Periods in a holomorphic family satisfy Picard–Fuchs equations arising from the Gauss–Manin connection (see [15, 14, 9]). For a one-parameter family of genus $g$ curves, the classical reduction typically yields a non-canonical scalar equation of order $2g$ [14]. For a generic family (Definition 12.1) we instead construct canonically a second-order equation with $g\times g$ matrix coefficients on the Hodge bundle, whose solutions are period vectors; the associated matrix Schwarzian produces characteristic invariants generalizing Dedekind’s genus-one computation.
  2. (ii)
    Periods of higher-dimensional varieties: cubic threefolds (Section 14). For a one-parameter family of smooth cubic threefolds in $\mathbb{P}^{4}$, the Gauss–Manin connection gives a first-order system on $H^{3}$ of rank $10$. Our approach produces a canonical second-order system on the rank-55 Hodge bundle $E=H^{2,1}$, yielding Schwarzian-type invariants for this higher-dimensional variation of Hodge structure.
  3. (iii)
    Mass–spring systems (Section 15). Mass–spring systems naturally give rise to second-order ordinary differential equations with matrix-valued coefficients. This example clarifies the philosophy of the paper: time is treated as a curve and the coefficients become geometric data along that curve, transforming coherently under changes of local time coordinate.
Organization of the paper. The paper is divided into five parts.
Part I (Sections 2–3) recalls the classical Schwarzian invariant of a second-order scalar ODE and interprets it as curvature.
Part II (Sections 4–8) develops a (quantum) theory of connections and differentials on Riemann surfaces.
Part III (Section 11) discusses modular forms and modular connections.
Part IV (Sections 12–15) treats periods of Riemann surfaces and cubic threefolds from a Schwarzian perspective.
Part V (Sections 16–19) discusses the classical and quantum damped mass–spring system in mechanics and its relation to the Schwarzian.
Part I Classical Schwarzian Derivative
2. Second order scalar ODE
Let $U\subset\mathbb{C}$ be a domain with coordinate $z$, and let $p,q\in\mathcal{M}(U)$ be meromorphic functions. We consider the second order linear equation
(2.1) $y^{\prime\prime}=2p\,y^{\prime}+q\,y.$
If $y_{1},y_{2}$ are two independent solutions, then the ratio $y_{1}/y_{2}$ is locally univalent and is well defined only up to a Möbius transformation (change of basis in the two-dimensional solution space). The Schwarzian derivative,
$S(w):=\frac{w^{\prime\prime\prime}}{w^{\prime}}-\frac{3}{2}\Bigl(\frac{w^{\prime\prime}}{w^{\prime}}\Bigr)^{2}=\Bigl(\frac{w^{\prime\prime}}{w^{\prime}}\Bigr)^{\prime}-\frac{1}{2}\Bigl(\frac{w^{\prime\prime}}{w^{\prime}}\Bigr)^{2},$
is invariant under Möbius transformations, hence $S(y_{1}/y_{2})$ is canonically attached to the equation.
The following classical computation is standard (see, for example, [11, § IV.2]). For a classical discussion of the projective viewpoint on second-order equations, see also Klein’s Lectures on the Icosahedron [16, Chapter III, §7].
Proposition 2.1.
Let $y_{1},y_{2}$ be two linearly independent solutions of (2.1). Then the Schwarzian derivative $S(y_{1}/y_{2})$ is independent of the chosen ordered basis of solutions and satisfies
(2.2) $S\Bigl(\frac{y_{1}}{y_{2}}\Bigr)=2\bigl(p^{\prime}-p^{2}-q\bigr).$
Comment. The right-hand side $2(p^{\prime}-p^{2}-q)$ is invariant under gauge changes of the dependent variable $y\mapsto f\,y$ (with $f$ nowhere vanishing), and under a change of coordinate it transforms with the usual Schwarzian anomaly. This is precisely the “projective curvature” viewpoint that motivates the non-abelian generalizations developed later.
The identity (2.2) implies that $2(p^{\prime}-p^{2}-q)$ is invariant under the two basic operations on (2.1):
  • (i)
    Gauge (change of dependent variable). If $y=f\,\widetilde{y}$ with $f$ nowhere vanishing, then $\widetilde{y}$ satisfies $\widetilde{y}^{\prime\prime}=2\widetilde{p}\,\widetilde{y}^{\prime}+\widetilde{q}\,\widetilde{y}$ with
    $\widetilde{p}=p-\frac{f^{\prime}}{f},\qquad\widetilde{q}=q+2p\frac{f^{\prime}}{f}-\frac{f^{\prime\prime}}{f},$
[…]
  • (ii)
    Change of coordinate. If $z=\lambda(\zeta)$ is a biholomorphism and $\widetilde{y}(\zeta)=y(\lambda(\zeta))$, then $\widetilde{y}$ satisfies $\widetilde{y}^{\prime\prime}=2\widetilde{p}\,\widetilde{y}^{\prime}+\widetilde{q}\,\widetilde{y}$ with
    $\widetilde{p}=(p\circ\lambda)\lambda^{\prime}+\frac{1}{2}\frac{\lambda^{\prime\prime}}{\lambda^{\prime}},\qquad\widetilde{q}=(q\circ\lambda)(\lambda^{\prime})^{2}.$
    In particular,
[…]
Thus $2(p^{\prime}-p^{2}-q)$ transforms as a projective connection.
3. Projective curvature and the Schwarzian anomaly
Equation (2.1) suggests that the data of $(p,q)$ should be organized so that the projective curvature (2.2) becomes intrinsic.
Connections of eccentricity 1/2.
The transformation law in (ii) shows that $p$ behaves like a connection coefficient: under $z=\lambda(\zeta)$ one has
$\widetilde{p}=(p\circ\lambda)\lambda^{\prime}+\frac{1}{2}\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}.$
This is the one-dimensional avatar of the usual gauge transformation law for a connection, except that the change of coordinate contributes an inhomogeneous term. The corresponding (curvature-type) combination is $p^{\prime}-p^{2}$.
Projective anomaly.
The expression
$r:=2(p^{\prime}-p^{2}-q)$
is the classical projective curvature attached to (2.1); it equals $S(y_{1}/y_{2})$ by Proposition 2.1. Under coordinate change $z=\lambda(\zeta)$ it satisfies
$\widetilde{r}=(r\circ\lambda)(\lambda^{\prime})^{2}+S(\lambda),$
so $r$ is not a quadratic differential in general, but it fails to be one by a scalar term given by the Schwarzian derivative of the transition map. In particular, if all coordinate changes are Möbius transformations (so $S(\lambda)=0$), then $r$ is a genuine quadratic differential. This is the mechanism behind the special features of the modular case.
Schwarzian as curvature of a connection.
Let $f$ be a locally univalent meromorphic function on $U\subset\mathbb{C}$. On the open set where $f^{\prime}\neq 0$ the logarithmic derivative
$A(f)=\frac{f^{\prime\prime}}{f^{\prime}}$
behaves like a connection: for any biholomorphism $\lambda$ one has the coordinate-change rule
$A(f\circ\lambda)=\lambda^{\prime}\bigl(A(f)\circ\lambda\bigr)+\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}.$
The curvature-type expression
$F_{A}(f)=\left(\frac{f^{\prime\prime}}{f^{\prime}}\right)^{\prime}-\frac{1}{2}\left(\frac{f^{\prime\prime}}{f^{\prime}}\right)^{2}$
is exactly the Schwarzian derivative $S(f)$. The projective anomaly is the Schwarzian chain rule
$S(f\circ\lambda)=(\lambda^{\prime})^{2}\bigl(S(f)\circ\lambda\bigr)+S(\lambda),$
and $S$ is invariant under postcomposition of $f$ by a Möbius transformation. This point of view motivates the non-abelian constructions developed in the next part.
Part II Quantum vs classical
4. Classical differentials and connections
Let $X$ be a Riemann surface. Choose a covering by coordinate charts $\{U_{i}\}_{i}$ on $X$. By abuse of notation we identify each $U_{i}$ with an open subset of $\mathbb{C}$. For each pair $i,j$ with $U_{i}\cap U_{j}\neq\varnothing$, the change of coordinate induces a biholomorphism
$\lambda_{ij}:U_{i}\cap U_{j}\longrightarrow U_{i}\cap U_{j},$
where the domain is viewed as an open subset of $\mathbb{C}$ via the coordinate on $U_{j}$, and the codomain is viewed as an open subset of $\mathbb{C}$ via the coordinate on $U_{i}$. The maps $\lambda_{ij}$ satisfy the usual cocycle conditions.
Recall that $K_{X}$ denotes the canonical line bundle of holomorphic 11-forms on $X$. Fix an integer $m$. A meromorphic connection on $K_{X}^{m}$ may be encoded locally by meromorphic functions $a_{i}\in\mathcal{M}(U_{i})$ satisfying on overlaps
$a_{j}=\lambda^{\prime}_{ij}(a_{i}\circ\lambda_{ij})-m\frac{\lambda^{\prime\prime}_{ij}}{\lambda^{\prime}_{ij}}.$
A classical meromorphic $m$-differential on $X$ is a meromorphic section of $K_{X}^{m}$. Equivalently, it is given by meromorphic functions $\omega_{i}\in\mathcal{M}(U_{i})$ satisfying on overlaps
$\omega_{j}=(\lambda^{\prime}_{ij})^{m}\,(\omega_{i}\circ\lambda_{ij}).$
More generally, if $V$ is a finite-dimensional complex vector space, a $V$-valued classical $m$-differential is a meromorphic section of $K_{X}^{m}\otimes V$, and in local coordinates it is represented by meromorphic maps $\omega_{i}:U_{i}\to V$ with the same tensorial rule.
Similarly, a meromorphic connection on $K_{X}^{m}$ may be encoded by meromorphic functions $a_{i}\in\mathcal{M}(U_{i})$ satisfying
$a_{j}=\lambda^{\prime}_{ij}(a_{i}\circ\lambda_{ij})-m\frac{\lambda^{\prime\prime}_{ij}}{\lambda^{\prime}_{ij}}.$
If $e\in\mathbb{Z}$, this is the same as a connection on $K_{X}^{-e}$ with m=em=-e; the parameter $e$ is the eccentricity used later in the quantum transformation law.
When we allow coefficients in $\operatorname{End}(V)$, the classical objects above correspond exactly to the $\mathcal{O}$-linear quantum objects introduced in Section 5 below:
Example 4.1.
A quantum connection (resp. a quantum $m$-differential) with values in $V$ is called classical with values in $\operatorname{End}(V)$ if it is $\mathcal{O}_{X}$-linear. Equivalently, it is given by multiplication by a coefficient $a_{i}\in\mathcal{M}(U_{i},\operatorname{End}(V))$ (resp. $\psi_{i}\in\mathcal{M}(U_{i},\operatorname{End}(V))$):
$A_{i}(f)=a_{i}f,\qquad\Psi_{i}(f)=\psi_{i}f.$
The coefficients satisfy the corresponding change of coordinate rules. In particular, an $\mathcal{O}_{X}$-linear quantum connection of eccentricity $e$ corresponds to a classical connection on $K_{X}^{-e}$ with coefficients in $\operatorname{End}(V)$, and an $\mathcal{O}_{X}$-linear quantum $m$-differential corresponds to a classical section of $K_{X}^{m}\otimes\operatorname{End}(V)$.
This classical vs. quantum distinction is important throughout the paper: in many geometric situations the most natural operations (Maurer–Cartan connections, Wronskians…) are not $\mathcal{O}$-linear.
5. Quantum differentials and connections
We now pass from $\mathcal{O}$-linear tensors to operator-valued objects. The basic input is that many natural constructions (Maurer–Cartan, Wronskians, exponentials, etc.) are only defined on Zariski open subsets depending on the input function. The following definition packages this “generic” behavior in a coordinate-free way.
Many of the operators that appear naturally in this paper (for instance $f\mapsto f^{\prime\prime}/f^{\prime}$ or $f\mapsto e^{f}$) do not produce a meromorphic function on the entire chart $U_{i}$ even when the input $f$ is meromorphic on $U_{i}$. Rather, the output is meromorphic on a Zariski open subset obtained by deleting a discrete set of points depending on $f$ (for $f^{\prime\prime}/f^{\prime}$ one removes the zeros of $f^{\prime}$, while for $e^{f}$ one removes the poles of $f$).
To keep the formalism honest while retaining the coordinate-free cocycle laws, we systematically work “generically” on charts. If $U\subset X$ is open, we call $U^{\circ}\subset U$ Zariski open if $U\setminus U^{\circ}$ is a discrete subset. For a complex vector space $V$ we set
[…]
the direct limit with respect to restriction. Thus an element of $\mathcal{M}_{\mathrm{gen}}(U,V)$ is represented by a $V$-valued meromorphic function on some Zariski open subset of $U$, and two representatives are identified if they agree after restriction to a smaller Zariski open subset.
All identities below involving quantum connections and quantum differentials are understood in this “generic” sense: the equality of two expressions means that they agree on some Zariski open subset on which both sides are defined. In degenerate cases (for example $f^{\prime\prime}/f^{\prime}$ when $f$ is constant) the natural domain of definition may be empty; such exceptional inputs occur in very special situations and will be ignored.
Definition 5.1.
Fix $e\in\mathbb{C}$ (the eccentricity) and a complex vector space $V$. A quantum connection on $X$ of eccentricity $e$ with values in $V$ is a family of local operators
$A_{i}:\mathcal{M}_{\mathrm{gen}}(U_{i},V)\dashrightarrow\mathcal{M}_{\mathrm{gen}}(U_{i},V),$
such that on every overlap $U_{i}\cap U_{j}$ one has, as operators on $\mathcal{M}_{\mathrm{gen}}(U_{i}\cap U_{j},V)$,
$A_{j}=\lambda^{\prime}_{ij}(A_{i}\circ\lambda_{ij})+e\frac{\lambda^{\prime\prime}_{ij}}{\lambda^{\prime}_{ij}}I.$
Here II denotes the identity operator on $\mathcal{M}_{\mathrm{gen}}(U_{i}\cap U_{j},V)$, scalar functions act by pointwise multiplication, and the transported operator is
$(A_{i}\circ\lambda_{ij})(f):=\bigl(A_{i}(f\circ\lambda_{ij}^{-1})\bigr)\circ\lambda_{ij}.$
Remark 5.2.
In the scalar case, the Maurer–Cartan operator $A(f):=f^{\prime\prime}/f^{\prime}$ (defined on the Zariski open subset where $f^{\prime}\neq 0$) is a basic example of a quantum connection of eccentricity $e=1$. Indeed, for a change of coordinate $\lambda$ one has $(f\circ\lambda)^{\prime}=(f^{\prime}\circ\lambda)\lambda^{\prime}$ and
$\frac{(f\circ\lambda)^{\prime\prime}}{(f\circ\lambda)^{\prime}}=\lambda^{\prime}\left(\frac{f^{\prime\prime}}{f^{\prime}}\circ\lambda\right)+\frac{\lambda^{\prime\prime}}{\lambda^{\prime}},$
i.e. $A(f\circ\lambda)=\lambda^{\prime}\,(A(f)\circ\lambda)+\lambda^{\prime\prime}/\lambda^{\prime}$.
Definition 5.3.
Fix an integer $m$ and a complex vector space $V$. A quantum $m$-differential on $X$ with values in $V$ is a family of local operators
$\Psi_{i}:\mathcal{M}_{\mathrm{gen}}(U_{i},V)\dashrightarrow\mathcal{M}_{\mathrm{gen}}(U_{i},V),$
such that on every overlap $U_{i}\cap U_{j}$ one has, as operators on $\mathcal{M}_{\mathrm{gen}}(U_{i}\cap U_{j},V)$,
$\Psi_{j}=(\lambda^{\prime}_{ij})^{m}(\Psi_{i}\circ\lambda_{ij}),$
where
$(\Psi_{i}\circ\lambda_{ij})(f):=\bigl(\Psi_{i}(f\circ\lambda_{ij}^{-1})\bigr)\circ\lambda_{ij}.$
In the definition of a quantum $m$-differential one may more generally allow operators
$\Psi_{i}:\mathcal{M}_{\mathrm{gen}}(U_{i},V)\dashrightarrow\mathcal{M}_{\mathrm{gen}}(U_{i},W),$
for (possibly different) complex vector spaces $V$ and $W$; the same transformation rule applies.
Examples.
Example 5.4.
Let $\omega$ be an ordinary meromorphic $m$-differential on $X$. In local coordinates it is given by meromorphic functions $\omega_{i}$ satisfying $\omega_{j}=(\lambda^{\prime}_{ij})^{m}(\omega_{i}\circ\lambda_{ij})$. Then multiplication by $\omega$,
$\Psi_{i}(f):=\omega_{i}f,$
defines a quantum $m$-differential (with values in any $V$).
Example $5.$
Assume $m\geq 0$. Define locally $\Psi_{i}(f):=(f^{\prime})^{m}$. Then $\Psi$ is a quantum $m$-differential since $(f\circ\lambda)^{\prime}=(f^{\prime}\circ\lambda)\lambda^{\prime}$ implies $((f\circ\lambda)^{\prime})^{m}=(\lambda^{\prime})^{m}((f^{\prime})^{m}\circ\lambda)$.
Example 5.6 (Wronskian for vector-valued functions).
Let $V$ be a complex vector space of dimension $n$. For $f\in\mathcal{M}(U_{i},V)$ define
$W_{i}(f):=f\wedge f^{\prime}\wedge f^{\prime\prime}\wedge\cdots\wedge f^{(n-1)}\in\bigwedge^{n}V.$
Then $W=\{W_{i}\}$ is a quantum $\frac{n(n-1)}{2}$-differential with values in $\bigwedge^{n}V$.
Example 5.7 (Wronskian of a quantum differential).
Let $r\in\mathbb{Z}$ and let $V$ be a complex vector space of dimension $n$. Suppose $\Psi$ is a quantum $r$-differential with values in $V$. For $k\geq 0$ define $\Psi^{(k)}$ locally by $\Psi^{(k)}_{i}(f):=(\Psi_{i}(f))^{(k)}$. Note that $\Psi^{(1)}$ need not be a quantum differential. Nevertheless the Wronskian
$\operatorname{Wr}(\Psi)_{i}(f):=\Psi_{i}(f)\wedge\Psi^{(1)}_{i}(f)\wedge\cdots\wedge\Psi^{(n-1)}_{i}(f)\in\bigwedge^{n}V$
defines a quantum $\bigl(nr+\frac{n(n-1)}{2}\bigr)$-differential with values in $\bigwedge^{n}V$.
Lemma 5.8 (Left Maurer–Cartan connection).
Let $V$ be a finite-dimensional complex vector space and let $\Psi=\{\Psi_{i}\}$ be a quantum $m$-differential on $X$ with values in $\operatorname{End}(V)$, where $m\neq 0$. For each $i$ define
$A_{i}(f):=\frac{1}{m}\Psi_{i}(f)^{-1}\Psi^{\prime}_{i}(f)$
on the open subset of $U_{i}$ where $\Psi_{i}(f)$ is pointwise invertible. Then $A=\{A_{i}\}$ is a quantum connection of eccentricity 11 with values in $\operatorname{End}(V)$.
Proof.
Differentiate the relation $\Psi_{j}=(\lambda^{\prime})^{m}(\Psi_{i}\circ\lambda)$ and multiply on the left by $\Psi_{j}^{-1}$. A direct computation shows that the logarithmic derivative $\frac{1}{m}\Psi^{-1}\Psi^{\prime}$ transforms as $A_{j}=\lambda^{\prime}(A_{i}\circ\lambda)+\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}I$. ∎
Example 5.9 (A scalar connection from a Wronskian).
Let $m\in\mathbb{Z}$ and let $V$ be a complex vector space of dimension $n$. If $\Psi$ is a quantum $m$-differential with values in $V$, then the previous example shows that $\operatorname{Wr}(\Psi)$ is a quantum
$N:=mn+\frac{n(n-1)}{2}=\frac{n(n+2m-1)}{2}$
-differential with values in the one-dimensional space $\bigwedge^{n}V$. On the open subset where $\operatorname{Wr}(\Psi)(f)\neq 0$ one may form the scalar quotient
$\frac{\bigl(\operatorname{Wr}(\Psi)_{i}(f)\bigr)^{\prime}}{\operatorname{Wr}(\Psi)_{i}(f)}.$
Defining
$A_{i}(f):=\frac{1}{N}\frac{\bigl(\operatorname{Wr}(\Psi)_{i}(f)\bigr)^{\prime}}{\operatorname{Wr}(\Psi)_{i}(f)}=\frac{2}{n(n+2m-1)}\frac{\bigl(\operatorname{Wr}(\Psi)_{i}(f)\bigr)^{\prime}}{\operatorname{Wr}(\Psi)_{i}(f)},$
one obtains a scalar quantum connection of eccentricity 11.
Example 5.10 (Classical connections and $m$-differentials with values in $\operatorname{End}(V)$).
A quantum connection (resp. a quantum $m$-differential) with values in $V$ is called classical with values in $\operatorname{End}(V)$ if it is $\mathcal{O}_{X}$-linear. Equivalently, it is given by multiplication by a coefficient $a_{i}\in\mathcal{M}(U_{i},\operatorname{End}(V))$ (resp. $\psi_{i}\in\mathcal{M}(U_{i},\operatorname{End}(V))$):
$A_{i}(f)=a_{i}f,\qquad\Psi_{i}(f)=\psi_{i}f.$
The coefficients satisfy the corresponding change of coordinate rules. In particular, an $\mathcal{O}_{X}$-linear quantum connection of eccentricity $e$ corresponds to a classical connection on $K_{X}^{-e}$ with coefficients in $\operatorname{End}(V)$, and an $\mathcal{O}_{X}$-linear quantum $m$-differential corresponds to a classical section of $K_{X}^{m}\otimes\operatorname{End}(V)$.
6. Curvature
The classical identity $S(f)=(f^{\prime\prime}/f^{\prime})^{\prime}-\tfrac{1}{2}(f^{\prime\prime}/f^{\prime})^{2}$ suggests that the correct “curvature” of an operator-valued connection should involve both a derivative term and a quadratic term. In the quantum setting we also have to keep track of the coordinate-change anomaly.
Definition 6.1.
Let AA be a quantum connection on $X$ with values in $\operatorname{End}(V)$ and eccentricity $e\neq 0$. For $m\in\mathbb{Z}$ and a quantum $m$-differential $\Psi$ with values in $V$, define the quantum covariant derivative $\nabla_{A}\Psi$ locally by
$(\nabla_{A}\Psi)_{i}(f):=\Psi_{i}(f)^{\prime}-\frac{m}{e}A_{i}(f)\Psi_{i}(f).$
Proposition 6.2.
If AA has eccentricity $e\neq 0$ and $\Psi$ is a quantum $m$-differential, then $\nabla_{A}\Psi$ is a quantum $(m+1)$-differential.
Proof.
On an overlap with change of coordinate $\lambda$, differentiate the identity $\Psi_{j}=(\lambda^{\prime})^{m}(\Psi_{i}\circ\lambda)$ and use the transformation rule $A_{j}=\lambda^{\prime}(A_{i}\circ\lambda)+e\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}I$. The coefficient $\frac{m}{e}$ is chosen precisely so that the $\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}$-terms cancel, yielding
$(\nabla_{A}\Psi)_{j}=(\lambda^{\prime})^{m+1}\bigl((\nabla_{A}\Psi)_{i}\circ\lambda\bigr).$
Remark 6.3 (Classical vs. quantum).
Even if AA and $\Psi$ are classical in the sense of the previous example, the operator $\nabla_{A}\Psi$ need not be classical: the derivative term $\Psi_{i}(f)^{\prime}$ typically produces an $f^{\prime}$-term. In the classical setting one instead defines the coefficient-level covariant derivative by
$(\nabla_{a}\psi)_{i}:=\psi^{\prime}_{i}-\frac{m}{e}a_{i}\psi_{i},$
which is again classical. Thus the classical and quantum covariant derivatives are distinct constructions.
We briefly recall the classical Cartan structural equation for a connection with values in a Lie algebra. Let $G$ be a Lie group with Lie algebra $\mathfrak{g}$. A $\mathfrak{g}$-valued connection 11-form on a manifold $M$ is a $\mathfrak{g}$-valued differential 11-form $A\in\Omega^{1}(M,\mathfrak{g})$. Its curvature is the $\mathfrak{g}$-valued 22-form
$F=dA+\frac{1}{2}[A,A],$
equivalently $F=dA+A\wedge A$. Under a gauge transformation $g:M\to G$ one has $A\mapsto g^{-1}Ag+g^{-1}dg$ and $F\mapsto g^{-1}Fg$.
Definition 6.4.
Let AA be a quantum connection on $X$ with values in $\operatorname{End}(V)$ and eccentricity $e\neq 0$. The quantum curvature of AA is the family of local operators
$(F_{A})_{i}:\mathcal{M}_{\mathrm{gen}}(U_{i},\operatorname{End}(V))\dashrightarrow\mathcal{M}_{\mathrm{gen}}(U_{i},\operatorname{End}(V))$
defined by
$(F_{A})_{i}(f):=A_{i}(f)^{\prime}-\frac{1}{2e}A_{i}(f)^{2}.$
Lemma 6.5.
Let AA be a quantum connection of eccentricity $e\neq 0$ with values in $\operatorname{End}(V)$. On an overlap $U_{i}\cap U_{j}$ with change of coordinate $\lambda=\lambda_{ij}$, one has
$(F_{A})_{j}=(\lambda^{\prime})^{2}\bigl((F_{A})_{i}\circ\lambda\bigr)+e\,S(\lambda)I,$
where $S(\lambda)$ is the Schwarzian derivative of $\lambda$, and II is the identity endomorphism.
Proof.
By the defining transformation rule for a quantum connection,
$A_{j}\;=\;\lambda^{\prime}\,(A_{i}\circ\lambda)\;+\;e\,\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}\,I.$
Differentiating (with respect to the local coordinate on $U_{j}$) gives
$A_{j}^{\prime}\;=\;\lambda^{\prime\prime}(A_{i}\circ\lambda)\;+\;(\lambda^{\prime})^{2}(A_{i}^{\prime}\circ\lambda)\;+\;e\Bigl(\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}\Bigr)^{\prime}I.$
Moreover, since $\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}I$ is scalar, it commutes with everything, and
$A_{j}^{2}=(\lambda^{\prime})^{2}(A_{i}^{2}\circ\lambda)+2e\,\lambda^{\prime\prime}(A_{i}\circ\lambda)+e^{2}\Bigl(\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}\Bigr)^{2}I.$
Substituting into the definition $(F_{A})_{j}=A_{j}^{\prime}-\frac{1}{2e}A_{j}^{2}$, the $\lambda^{\prime\prime}(A_{i}\circ\lambda)$ terms cancel, yielding
$(F_{A})_{j}=(\lambda^{\prime})^{2}\Bigl(A_{i}^{\prime}-\frac{1}{2e}A_{i}^{2}\Bigr)\circ\lambda\;+\;e\left[\Bigl(\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}\Bigr)^{\prime}-\frac{1}{2}\Bigl(\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}\Bigr)^{2}\right]I.$
The bracketed expression is exactly $S(\lambda)$, so the claim follows. ∎
Remark 6.6 (Schwarzian as curvature).
Consider the scalar quantum 11-differential $\Psi(f)=f^{\prime}$. On the open set where $f^{\prime}\neq 0$, the Maurer–Cartan construction yields the scalar quantum connection $A(f)=f^{\prime\prime}/f^{\prime}$. Its quantum curvature is
$F_{A}(f)=\left(\frac{f^{\prime\prime}}{f^{\prime}}\right)^{\prime}-\frac{1}{2}\left(\frac{f^{\prime\prime}}{f^{\prime}}\right)^{2}=S(f).$
Applying the previous lemma gives the chain rule
$S(f\circ\lambda)=(\lambda^{\prime})^{2}\bigl(S(f)\circ\lambda\bigr)+S(\lambda).$
Since $S\!\left(\frac{az+b}{cz+d}\right)=0$ for every Möbius transformation ($ad-bc\neq 0$), it follows that $S\!\left(\frac{af+b}{cf+d}\right)=S(f)$ for every nonconstant meromorphic function $f$.
Gauge action.
Definition 6.7 (Gauge group).
Let $V$ be a finite-dimensional complex vector space. The gauge group $G$ consists of all classical invertible 0-differentials with values in $\operatorname{End}(V)$, i.e. meromorphic functions $g\in\mathcal{M}(X,\operatorname{End}(V))$ whose determinant is not identically zero.
Definition 6.8 (Gauge action).
Let AA be a quantum connection with values in $\operatorname{End}(V)$ and eccentricity $e$. For $g\in G$ define
$g\bullet A:=gAg^{-1}+g^{\prime}g^{-1},$
i.e. locally $(g\bullet A)_{i}=g_{i}A_{i}g_{i}^{-1}+g^{\prime}_{i}g_{i}^{-1}$.
We recall that the appearance of an affine “connection term” in the transformation law is classical; see Klein [16, Chapter III, §7] for the corresponding discussion in the genus one setting.
Proposition 6.9.
If AA is a quantum connection of eccentricity $e$, then $g\bullet A$ is again a quantum connection of eccentricity $e$. Moreover, $(g,A)\mapsto g\bullet A$ defines an action of $G$.
In particular, the curvature defined below is gauge covariant, while under change of coordinates it transforms with a Schwarzian projective anomaly.
Vector bundles on Riemann surfaces
Let $E\to X$ be a holomorphic vector bundle of rank $n$. Choose, for each $i$, a holomorphic frame of $E$ over $U_{i}$, so that $E|_{U_{i}}\simeq U_{i}\times\mathbb{C}^{n}$. On an overlap $U_{i}\cap U_{j}$ the corresponding transition function is a holomorphic map
$g_{ij}:U_{i}\cap U_{j}\longrightarrow\operatorname{GL}_{n}(\mathbb{C}),$
and a meromorphic section $s$ of $E$ is represented in these frames by meromorphic functions $s_{i}\in\mathcal{M}(U_{i},\mathbb{C}^{n})$ satisfying
$s_{j}=g_{ji}\,(s_{i}\circ\lambda_{ij})\qquad\text{on }U_{i}\cap U_{j},$
where $g_{ji}=g_{ij}^{-1}$ and the composition with $\lambda_{ij}$ expresses the same point in $U_{i}$-coordinates.
Recall the gauge action notation from Definition 6.8: for an invertible matrix-valued meromorphic function $g$ on $U_{i}\cap U_{j}$ we write
$g\bullet B:=g\,B\,g^{-1}+g^{\prime}g^{-1},$
for the usual gauge action on local operators (where $g$ acts by pointwise multiplication).
A quantum connection on $E$ of eccentricity $e\in\mathbb{C}$ is a family of local operators
$A_{i}:\mathcal{M}_{\mathrm{gen}}(U_{i},\mathbb{C}^{n})\dashrightarrow\mathcal{M}_{\mathrm{gen}}(U_{i},\mathbb{C}^{n})$
such that on every overlap $U_{i}\cap U_{j}$ one has, as operators on $\mathcal{M}_{\mathrm{gen}}(U_{i}\cap U_{j},\mathbb{C}^{n})$,
$A_{j}=g_{ji}\bullet\Bigl(\lambda^{\prime}_{ij}(A_{i}\circ\lambda_{ij})+e\frac{\lambda^{\prime\prime}_{ij}}{\lambda^{\prime}_{ij}}I\Bigr).$
Equivalently,
$A_{j}=g_{ji}\Bigl(\lambda^{\prime}_{ij}(A_{i}\circ\lambda_{ij})+e\frac{\lambda^{\prime\prime}_{ij}}{\lambda^{\prime}_{ij}}I\Bigr)g_{ij}\;+\;(g_{ji})^{\prime}g_{ij}.$
Here II is the identity operator, the matrices $g_{ij}$ act by pointwise multiplication, and primes denote differentiation with respect to the local coordinate on $U_{j}$.
Similarly, for $m\in\mathbb{Z}$, a quantum $m$-differential on $E$ is a family of local operators
$\Psi_{i}:\mathcal{M}_{\mathrm{gen}}(U_{i},\mathbb{C}^{n})\dashrightarrow\mathcal{M}_{\mathrm{gen}}(U_{i},\mathbb{C}^{n})$
such that on overlaps
$\Psi_{j}=(\lambda^{\prime}_{ij})^{m}\,g_{ji}\,(\Psi_{i}\circ\lambda_{ij})\,g_{ij}.$
More generally, if $E$ and $F$ are holomorphic vector bundles with transition functions $g^{E}_{ij}$ and $g^{F}_{ij}$, one may allow operators $\Psi_{i}:\mathcal{M}_{\mathrm{gen}}(U_{i},\mathbb{C}^{\mathrm{rk}\,E})\dashrightarrow\mathcal{M}_{\mathrm{gen}}(U_{i},\mathbb{C}^{\mathrm{rk}\,F})$ transforming by
$\Psi_{j}=(\lambda^{\prime}_{ij})^{m}\,g^{F}_{ji}\,(\Psi_{i}\circ\lambda_{ij})\,g^{E}_{ij}.$
In the trivial-bundle case $E=X\times V$ (so $g_{ij}=I$), these definitions reduce to Definitions 5.1 and 5.3.
Example (Multiplication by an $\operatorname{End}(E)$-valued $m$-differential).
Let $\omega$ be a classical meromorphic section of $K_{X}^{m}\otimes\operatorname{End}(E)$. In local coordinates and frames it is given by meromorphic matrices $\omega_{i}$ satisfying
$\omega_{j}=(\lambda^{\prime}_{ij})^{m}\,g_{ji}\,(\omega_{i}\circ\lambda_{ij})\,g_{ij}.$
Then multiplication by $\omega$,
$\Psi_{i}(f):=\omega_{i}f,$
defines a (classical) quantum $m$-differential on $E$.
Example (Classical connections on $E\otimes K_{X}^{-e}$).
Assume $e\in\mathbb{Z}$ and suppose $E\otimes K_{X}^{-e}$ is equipped with a meromorphic connection. In local coordinates and frames, it is represented by meromorphic matrices $a_{i}$ whose transformation rule is precisely the coefficient version of the above formula for $A_{j}$. Equivalently, the $\mathcal{O}_{X}$-linear operators $A_{i}(f):=a_{i}f$ form a bundle-valued quantum connection on $E$ of eccentricity $e$.
The notion of equivalence of two connections.
Definition 6.10.
Let $A_{1}$ and $A_{2}$ be quantum connections on $X$ with the same eccentricity $e\neq 0$ and values in $\operatorname{End}(V)$. We say that $A_{1}$ and $A_{2}$ are classically equivalent if there exists $g\in G$ such that
$F_{A_{2}}=gF_{A_{1}}g^{-1}.$
Remark 6.11.
Although $F_{A}$ is not globally tensorial, it transforms on overlaps by
$(F_{A})_{j}=(\lambda^{\prime})^{2}\bigl((F_{A})_{i}\circ\lambda\bigr)+eS(\lambda)I.$
The anomaly term $eS(\lambda)I$ is scalar (proportional to II). Since elements of $G$ act by conjugation and scalar endomorphisms commute with all elements of $\operatorname{End}(V)$, the relation $F_{A_{2}}=gF_{A_{1}}g^{-1}$ is independent of the choice of local coordinates. Thus the definition is well defined.
7. Second order quantum ODE
We discuss second order quantum differential equations on a Riemann surface. The starting point is the local scalar equation on a domain $U\subset\mathbb{C}$,
$y^{\prime\prime}=p\,y^{\prime}+q\,y,$
whose coefficients depend on the chosen coordinate. In the quantum setting we take the solution to be a quantum 0-differential $\Psi$, and we allow the coefficients to be quantum tensors (connections and 22-differentials) so that the notion becomes intrinsic on $X$.
More precisely, on each chart $U_{i}$ and for each meromorphic function $f$ on $U_{i}$ we obtain a meromorphic function $\Psi_{i}(f)$ (defined generically on a Zariski open subset) which satisfies a second order ODE whose coefficients depend on $f$:
$\Psi_{i}(f)^{\prime\prime}=p_{i}(f)\,\Psi_{i}(f)^{\prime}+q_{i}(f)\,\Psi_{i}(f).$
We first determine the transformation law of the coefficient operators $p_{i}$ and $q_{i}$ under change of coordinates.
Lemma 7.1.
Let $\Psi$ be a quantum 0-differential and suppose that on each chart $U_{i}$ it satisfies the local equation
$\Psi_{i}(f)^{\prime\prime}=p_{i}(f)\,\Psi_{i}(f)^{\prime}+q_{i}(f)\,\Psi_{i}(f),\qquad f\in\mathcal{M}_{\mathrm{gen}}(U_{i}),$
for some operators $p_{i}$ and $q_{i}$. Then on an overlap $U_{i}\cap U_{j}$ with change of coordinate $\lambda=\lambda_{ij}$, the coefficients transform according to
$p_{j}=\lambda^{\prime}\,(p_{i}\circ\lambda)+\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}\,I,\qquad q_{j}=(\lambda^{\prime})^{2}\,(q_{i}\circ\lambda),$
as operators on $\mathcal{M}_{\mathrm{gen}}(U_{i}\cap U_{j})$. In particular, $q$ is a quantum 22-differential, while $p$ transforms as a quantum connection of eccentricity 11.
Proof.
All computations are performed on a Zariski open subset of $U_{i}\cap U_{j}$ on which the relevant expressions are defined. Since $\Psi$ is a quantum 0-differential, on $U_{i}\cap U_{j}$ we have
$\Psi_{j}=\Psi_{i}\circ\lambda,\qquad\Psi_{j}(f)=\bigl(\Psi_{i}(f\circ\lambda^{-1})\bigr)\circ\lambda.$
Fix $f\in\mathcal{M}(U_{j})$ and set $g=f\circ\lambda^{-1}\in\mathcal{M}(U_{i})$ and $Y:=\Psi_{i}(g)$. Then
$\Psi_{j}(f)=Y\circ\lambda.$
Differentiating with respect to the $U_{j}$-coordinate gives
$\Psi_{j}(f)^{\prime}=(Y^{\prime}\circ\lambda)\lambda^{\prime},\qquad\Psi_{j}(f)^{\prime\prime}=(Y^{\prime\prime}\circ\lambda)(\lambda^{\prime})^{2}+(Y^{\prime}\circ\lambda)\lambda^{\prime\prime}.$
Using the equation on $U_{i}$,
$Y^{\prime\prime}=p_{i}(g)Y^{\prime}+q_{i}(g)Y,$
we obtain
$\Psi_{j}(f)^{\prime\prime}$ $=(\lambda^{\prime})^{2}\,(p_{i}(g)\circ\lambda)(Y^{\prime}\circ\lambda)+(\lambda^{\prime})^{2}\,(q_{i}(g)\circ\lambda)(Y\circ\lambda)+(Y^{\prime}\circ\lambda)\lambda^{\prime\prime}.$
Since $Y\circ\lambda=\Psi_{j}(f)$ and $Y^{\prime}\circ\lambda=\Psi_{j}(f)^{\prime}/\lambda^{\prime}$, we conclude
$\Psi_{j}(f)^{\prime\prime}=\Bigl(\lambda^{\prime}(p_{i}(g)\circ\lambda)+\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}\Bigr)\Psi_{j}(f)^{\prime}+(\lambda^{\prime})^{2}(q_{i}(g)\circ\lambda)\,\Psi_{j}(f).$
Replacing $g$ by $f\circ\lambda^{-1}$ yields the stated operator transformation laws. ∎
Remark 7.2.
By Lemma 7.1, the coefficient $p$ transforms as a quantum connection of eccentricity 11. Hence we may view $p$ itself as a connection and form the corresponding covariant derivative $\nabla_{p}$ acting on quantum differentials.
Since $\Psi$ is a quantum 0-differential, one has
$(\nabla_{p}\Psi)(f)=\Psi(f)^{\prime},$
and therefore
$(\nabla_{p}^{2}\Psi)(f)=(\nabla_{p}(\nabla_{p}\Psi))(f)=\Psi(f)^{\prime\prime}-p(f)\Psi(f)^{\prime}.$
Consequently, the local equation
$\Psi_{i}(f)^{\prime\prime}=p_{i}(f)\Psi_{i}(f)^{\prime}+q_{i}(f)\Psi_{i}(f)$
may be rewritten intrinsically as
$\nabla_{p}^{2}\Psi=q\,\Psi,$
which is independent of the choice of local coordinate.
Example 7.3 (A basic quantum second-order equation).
Let $U_{i}$ be a chart with coordinate $z$, and let
$A(f)=\frac{f^{\prime\prime}}{f^{\prime}}$
be the scalar quantum connection (of eccentricity 11) on $U_{i}$. Let
$B(f)=(f^{\prime})^{2}$
be the quantum 22-differential given by squaring the first derivative. Fix a constant $c\in\mathbb{C}$, and set
$p=A,\qquad q=\omega^{2}B.$
Consider the quantum second-order equation
$\Psi_{i}(f)^{\prime\prime}=p_{i}(f)\,\Psi_{i}(f)^{\prime}+q_{i}(f)\,\Psi_{i}(f),\qquad f\in\mathcal{M}(U_{i}).$
Writing $Y(z)=\Psi_{i}(f)(z)$, the equation becomes
$Y^{\prime\prime}-\frac{f^{\prime\prime}}{f^{\prime}}Y^{\prime}-\omega^{2}(f^{\prime})^{2}Y=0.$
Set $t=f(z)$ and write $Y(z)=\widetilde{Y}(t)$. Then
$Y^{\prime}=f^{\prime}\widetilde{Y}^{\prime},\qquad Y^{\prime\prime}=f^{\prime\prime}\widetilde{Y}^{\prime}+(f^{\prime})^{2}\widetilde{Y}^{\prime\prime}$,
so the equation reduces to
$(f^{\prime})^{2}\bigl(\widetilde{Y}^{\prime\prime}-\omega^{2}\widetilde{Y}\bigr)=0.$
Since $f^{\prime}\neq 0$ on the domain of definition, we obtain the constant–coefficient ODE
$\widetilde{Y}^{\prime\prime}=\omega^{2}\widetilde{Y}$.
Thus the general local solution is
$\widetilde{Y}(t)=C_{1}e^{\omega t}+C_{2}e^{-\omega t},$
and hence
$\Psi_{i}(f)(z)=C_{1}e^{\omega f(z)}+C_{2}e^{-\omega f(z)}.$
In Part V we interpret the above differential equation as a (quantum) mass–spring system with frequency $\omega$.
Gauge invariance
Unfortunately the quantum curvature $F_{A}$ of a quantum connection is not gauge invariant. This defect becomes particularly transparent when the eccentricity is $e=\tfrac{1}{2}$, since in that case we may consider the second order quantum ODE
$\Psi^{\prime\prime}=2A\Psi^{\prime}+q\Psi,$
where $\Psi$ is a quantum 0-differential and $q$ is a quantum 22-differential. Equivalently, since for $e=\tfrac{1}{2}$ and $m=0$ one has
$(\nabla_{A}^{2}\Psi)(f)=\Psi(f)^{\prime\prime}-2A(f)\Psi(f)^{\prime},$
the equation may be written intrinsically as $\nabla_{A}^{2}\Psi=q\Psi$.
7.1. Main Lemma
We are now ready to state our main lemma. Suppose we have an invertible solution $g$ of this equation with values in $\operatorname{End}(V)$, then the corresponding gauge transform of AA has curvature $F_{A}-q$ up to conjugation.
Lemma 7.4.
Let AA be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. Let $g$ be an $\operatorname{End}(V)$-valued quantum 0-differential such that for every chart $U_{i}$ and every $f\in\mathcal{M}(U_{i})$ the meromorphic function $g_{i}(f)\in\mathcal{M}(U_{i},\operatorname{End}(V))$ satisfies $\det(g_{i}(f))\not\equiv 0$ on $U_{i}$. Assume that for every $i$ and every $f\in\mathcal{M}(U_{i})$ one has, on the open subset where $g_{i}(f)$ is pointwise invertible,
[…]
Consider the gauge-transformed connection $g^{-1}\bullet A$ (in the sense of Section 3). Explicitly, on $U_{i}$ and for $f\in\mathcal{M}(U_{i})$ one has
$(g^{-1}\bullet A)_{i}(f)\;=\;g_{i}(f)^{-1}A_{i}(f)g_{i}(f)\;-\;g_{i}(f)^{-1}g_{i}(f)^{\prime}.$
Then the quantum curvature satisfies, pointwise in $f$,
$F_{g^{-1}\bullet A,\,i}(f)\;=\;g_{i}(f)^{-1}\bigl(F_{A,\,i}(f)-q_{i}(f)\bigr)g_{i}(f),$
where for $e=\tfrac{1}{2}$ one has $F_{A,\,i}(f)=A_{i}(f)^{\prime}-A_{i}(f)^{2}$.
Proof.
Fix a chart $U_{i}$ and an input $f\in\mathcal{M}(U_{i})$. On the open subset where $g:=g_{i}(f)$ is pointwise invertible, write for brevity
$A:=A_{i}(f),\qquad q:=q_{i}(f),\qquad\widetilde{A}:=(g^{-1}\bullet A)_{i}(f)=g^{-1}Ag-g^{-1}g^{\prime}.$
Since $e=\tfrac{1}{2}$, the curvature is $F_{A}=A^{\prime}-A^{2}$, and hence
$F_{\widetilde{A}}=(\widetilde{A})^{\prime}-\widetilde{A}^{\,2}.$
We first expand the derivative term. Using $(g^{-1})^{\prime}=-g^{-1}g^{\prime}g^{-1}$, we obtain
[…]
$=-g^{-1}g^{\prime}g^{-1}Ag+g^{-1}A^{\prime}g+g^{-1}Ag^{\prime}+g^{-1}g^{\prime}g^{-1}g^{\prime}-g^{-1}g^{\prime\prime}.$
Next we expand the square:
$\widetilde{A}^{\,2}$ $=(g^{-1}Ag-g^{-1}g^{\prime})^{2}$
$=g^{-1}A^{2}g+g^{-1}g^{\prime}g^{-1}g^{\prime}-g^{-1}Ag^{\prime}-g^{-1}g^{\prime}g^{-1}Ag.$
Subtracting (3) from (2) gives
$F_{\widetilde{A}}=g^{-1}A^{\prime}g-g^{-1}A^{2}g-g^{-1}(g^{\prime\prime}-2Ag^{\prime}).$
By hypothesis, $g$ satisfies $g^{\prime\prime}=2Ag^{\prime}+qg$, so $(g^{\prime\prime}-2Ag^{\prime})=qg$ and therefore
$F_{\widetilde{A}}=g^{-1}(A^{\prime}-A^{2})g-g^{-1}qg=g^{-1}(F_{A}-q)g.$
This proves the claim. ∎
The quantum Schwarzian.
In the scalar case, Proposition 2.1 shows that the Schwarzian of the ratio of two independent solutions is
$S\Bigl(\frac{y_{1}}{y_{2}}\Bigr)=2\bigl(p^{\prime}-p^{2}-q\bigr)=2(F_{p}-q).$
This motivates the following definition in the non-abelian quantum setting.
Definition 7.5.
Let AA be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. For the quantum second-order equation
(7.0) $\Psi^{\prime\prime}=2A\Psi^{\prime}+q\Psi\qquad\bigl(\text{equivalently, }\nabla_{A}^{2}\Psi=q\Psi\bigr),$
we define its quantum Schwarzian to be the $\operatorname{End}(V)$-valued quantity
$\mathcal{S}_{A,q}:=2\,(F_{A}-q).$
Remark 7.6.
The quantum Schwarzian is the natural analogue of the classical invariant in Proposition 2.1. On an overlap $U_{i}\cap U_{j}$ with change of coordinate $\lambda=\lambda_{ij}$, the curvature of an eccentricity $e=\tfrac{1}{2}$ connection satisfies the transformation law
$F_{A,j}=(\lambda^{\prime})^{2}\,(F_{A,i}\circ\lambda)+\tfrac{1}{2}S(\lambda)\,I,$
while $q$ transforms tensorially as a quantum 22-differential. Hence
$\mathcal{S}_{A,q,j}=(\lambda^{\prime})^{2}\,(\mathcal{S}_{A,q,i}\circ\lambda)+S(\lambda)\,I.$
In particular, $\mathcal{S}_{A,q}$ has the same anomaly under coordinate change as the curvature (up to the normalizing factor 22).
Lemma 7.4 explains the gauge–theoretic meaning of $\mathcal{S}_{A,q}$. Namely, if $g$ is an invertible $\operatorname{End}(V)$-valued solution of (7.0), then
$2F_{g^{-1}\bullet A}=g^{-1}\,\mathcal{S}_{A,q}\,g.$
Thus the quantum Schwarzian is precisely the curvature of the gauge-transformed connection $g^{-1}\bullet A$, up to conjugation. In the scalar case ($V=\mathbb{C}$) this reduces to gauge invariance, and it is the direct quantum analogue of the fact that $S(y_{1}/y_{2})$ is independent of the choice of a basis of solutions.
8. The gauge and $\star$ actions
Let AA be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. The second-order equation (7.0) depends on the choice of the pair $(A,q)$, while its quantum Schwarzian $\mathcal{S}_{A,q}=2(F_{A}-q)$ is a more intrinsic quantity. We now describe a natural action of quantum 11-differentials on such pairs which preserves $F_{A}-q$ (and hence $\mathcal{S}_{A,q}$).
Definition 8.1.
Let $u$ be an $\operatorname{End}(V)$-valued quantum 11-differential on $X$. We define
$u\star(A,q):=(A+u,\;q-uA-Au-u^{2}+u^{\prime}),$
where the expression is understood chartwise: on $U_{i}$ and for $f\in\mathcal{M}_{\mathrm{gen}}(U_{i})$,
$(u\star(A,q))_{i}(f)=\bigl(A_{i}(f)+u_{i}(f),\;q_{i}(f)-u_{i}(f)A_{i}(f)-A_{i}(f)u_{i}(f)-u_{i}(f)^{2}+u_{i}(f)^{\prime}\bigr),$
and all identities are interpreted on a common Zariski open subset on which the relevant expressions are defined.
Lemma 8.2.
The operation $\star$ defines an action of the additive group of $\operatorname{End}(V)$-valued quantum 11-differentials on the set of pairs $(A,q)$ with $e$. Equivalently, for quantum 11-differentials $u$ and $v$ one has
$v\star(u\star(A,q))=(u+v)\star(A,q).$
Proof.
This is a direct computation using bilinearity of the products and linearity of the derivative. Writing $(A^{\prime},q^{\prime}):=u\star(A,q)$, we have $A^{\prime}=A+u$ and $q^{\prime}=q-uA-Au-u^{2}+u^{\prime}$. Hence
$v\star(A^{\prime},q^{\prime})$ $=\bigl(A+u+v,\;q-uA-Au-u^{2}+u^{\prime}-v(A+u)-(A+u)v-v^{2}+v^{\prime}\bigr)$
$=\bigl(A+u+v,\;q-(u+v)A-A(u+v)-(u+v)^{2}+(u+v)^{\prime}\bigr)$
$=(u+v)\star(A,q),$
as claimed. ∎
Lemma 8.3.
For $e=\tfrac{1}{2}$ the quantum Schwarzian is invariant under the $\star$-action. In other words,
$\mathcal{S}_{A,q}=\mathcal{S}_{u\star(A,q)}.$
Proof.
For $e=\tfrac{1}{2}$ one has $F_{A}=A^{\prime}-A^{2}$, hence
$F_{A+u}=(A+u)^{\prime}-(A+u)^{2}=F_{A}+u^{\prime}-Au-uA-u^{2}.$
Subtracting the transformed $q$ gives
$F_{A+u}-\bigl(q-uA-Au-u^{2}+u^{\prime}\bigr)=F_{A}-q,$
as desired. ∎
8.1. Main Theorem
Theorem 8.4 (Main Theorem).
Let $g\in G$ be a gauge transformation (Definition 6.7). Let AA be an $\operatorname{End}(V)$-valued quantum connection of eccentricity $e=\tfrac{1}{2}$ and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. Define
$A^{g}:=g^{-1}\bullet A=g^{-1}Ag-g^{-1}g^{\prime},\qquad q^{g}:=g^{-1}qg+2\,g^{-1}Ag^{\prime}-g^{-1}g^{\prime\prime}.$
Then $A^{g}$ is again a quantum connection of eccentricity $e=\tfrac{1}{2}$, and $q^{g}$ is an $\operatorname{End}(V)$-valued quantum 22-differential. Moreover, if $\Psi$ is a (quantum) 0-differential solving
$\Psi^{\prime\prime}=2A\Psi^{\prime}+q\Psi,$
then $\widetilde{\Psi}:=g^{-1}\Psi$ solves the transformed equation
$\widetilde{\Psi}^{\prime\prime}=2A^{g}\widetilde{\Psi}^{\prime}+q^{g}\widetilde{\Psi}$.
In particular,
$F_{A^{g}}-q^{g}=g^{-1}(F_{A}-q)g,\qquad\text{hence}\qquad\mathcal{S}_{A^{g},q^{g}}=g^{-1}\mathcal{S}_{A,q}g.$
Consequently the coefficients of the characteristic polynomial of $\mathcal{S}_{A,q}$ (equivalently, the functions $\operatorname{tr}(\mathcal{S}_{A,q}^{r})$) are invariants of the gauge-equivalence class of the equation. In particular, if $g$ itself is a (classical) $\operatorname{GL}(V)$-valued solution of $g^{\prime\prime}=2Ag^{\prime}+qg,$ then $q^{g}=0$ and $F_{A^{g}}=g^{-1}(F_{A}-q)g$, recovering Lemma 7.4.
Proof.
The first statement is immediate from the fact that the gauge action preserves the class of quantum connections (apply Proposition following Definition 6.8 to $g^{-1}$). For the transformation of the equation, write $\Psi=g\widetilde{\Psi}$. Then
$\Psi^{\prime}=g^{\prime}\widetilde{\Psi}+g\widetilde{\Psi}^{\prime},\qquad\Psi^{\prime\prime}=g^{\prime\prime}\widetilde{\Psi}+2g^{\prime}\widetilde{\Psi}^{\prime}+g\widetilde{\Psi}^{\prime\prime}$.
Substituting into $\Psi^{\prime\prime}=2A\Psi^{\prime}+q\Psi$ and multiplying on the left by $g^{-1}$ yields
$\widetilde{\Psi}^{\prime\prime}=2(g^{-1}Ag-g^{-1}g^{\prime})\,\widetilde{\Psi}^{\prime}+\bigl(g^{-1}qg+2g^{-1}Ag^{\prime}-g^{-1}g^{\prime\prime}\bigr)\,\widetilde{\Psi}$,
which is the claimed equation.
Finally, since $e=\tfrac{1}{2}$ we have $F_{A}=A^{\prime}-A^{2}$. A direct computation from $A^{g}=g^{-1}Ag-g^{-1}g^{\prime}$ gives
$F_{A^{g}}=g^{-1}F_{A}g+2\,g^{-1}Ag^{\prime}-g^{-1}g^{\prime\prime},$
and subtracting $q^{g}$ yields $F_{A^{g}}-q^{g}=g^{-1}(F_{A}-q)g$. Multiplying by 22 gives the stated transformation of $\mathcal{S}_{A,q}$. Conjugation invariance of characteristic polynomial coefficients and traces is immediate.
qedhere ∎
Remark 8.5.
It is important that the transformation of $q$ in Lemma 8.4 is not given by tensorial conjugation alone. In general the curvature $F_{A}=A^{\prime}-A^{2}$ is not gauge covariant under the action $A\mapsto g\bullet A=gAg^{-1}+g^{\prime}g^{-1}$ when AA is $\operatorname{End}(V)$-valued and noncommutative. The additional terms $2g^{-1}Ag^{\prime}-g^{-1}g^{\prime\prime}$ in $q^{g}$ are exactly what is needed so that the Schwarzian combination $F_{A}-q$ transforms by conjugation, and hence $\mathcal{S}_{A,q}=2(F_{A}-q)$ is the natural gauge-theoretic invariant of the second order equation.
Lemma 8.6 (Compatibility of the gauge and $\star$ actions).
Let AA be an $\operatorname{End}(V)$-valued quantum connection of eccentricity $e=\tfrac{1}{2}$ and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. Let $u$ be an $\operatorname{End}(V)$-valued quantum 11-differential and let $g\in G$ be a gauge transformation. Write
$u\star(A,q)=(A+u,\;\widetilde{q})\qquad\text{with}\qquad\widetilde{q}:=q-uA-Au-u^{2}+u^{\prime}.$
[…]
$\bigl((A+u)^{g},\widetilde{q}^{g}\bigr)\;=\;u^{g}\star(A^{g},q^{g}).$
Equivalently, the gauge group acts by automorphisms of the $\star$-action, so $\star$-orbits of pairs $(A,q)$ are preserved under gauge transformations.
Proof.
The first components agree:
$(A+u)^{g}=g^{-1}(A+u)g-g^{-1}g^{\prime}=(g^{-1}Ag-g^{-1}g^{\prime})+g^{-1}ug=A^{g}+u^{g}.$
For the second component, it is enough to compare the gauge-covariant combination $F_{A}-q$. By Lemma 8.3 one has $F_{A+u}-\widetilde{q}=F_{A}-q$, and by Lemma 8.4 this implies
$F_{(A+u)^{g}}-\widetilde{q}^{g}=g^{-1}(F_{A}-q)g.$
On the other hand, Lemma 8.4 gives $F_{A^{g}}-q^{g}=g^{-1}(F_{A}-q)g$, and applying Lemma 8.3 to the pair $(A^{g},q^{g})$ and the 11-differential $u^{g}$ yields
$F_{A^{g}+u^{g}}-q_{u^{g}\star(A^{g},q^{g})}=F_{A^{g}}-q^{g}=g^{-1}(F_{A}-q)g.$
Since $(A+u)^{g}=A^{g}+u^{g}$, the two pairs $\bigl((A+u)^{g},\widetilde{q}^{g}\bigr)$ and $u^{g}\star(A^{g},q^{g})$ have the same first component and the same value of $F_{A}-q$; therefore their second components coincide, proving the claim. ∎
Remark 8.7.
The action $\star$ is transitive on the choice of the connection component: if $A_{1}$ and $A_{2}$ are two eccentricity $e=\tfrac{1}{2}$ quantum connections, then $u:=A_{2}-A_{1}$ is a quantum 11-differential and $A_{2}=A_{1}+u$. Thus $\star$ parametrizes different presentations $(A,q)$ of the same quantum Schwarzian. In the classical scalar case this recovers the familiar fact that a projective connection can be represented by an affine connection modulo addition of a 11-form.
Example 8.8 (The quantum Schwarzian in Example 4.3).
Consider the scalar equation of Example 4.3,
$Y^{\prime\prime}=\frac{f^{\prime\prime}}{f^{\prime}}\,Y^{\prime}+c^{2}(f^{\prime})^{2}\,Y,$
and rewrite it in the eccentricity $e=\tfrac{1}{2}$ form
$Y^{\prime\prime}=2p\,Y^{\prime}+q\,Y$
by setting
$p=\frac{1}{2}\,\frac{f^{\prime\prime}}{f^{\prime}},\qquad q=c^{2}(f^{\prime})^{2}.$
Then the corresponding (scalar) quantum Schwarzian is
$\mathcal{S}_{p,q}=2\,(F_{p}-q)=2\bigl(p^{\prime}-p^{2}-q\bigr)=\left(\frac{f^{\prime\prime}}{f^{\prime}}\right)^{\prime}-\frac{1}{2}\left(\frac{f^{\prime\prime}}{f^{\prime}}\right)^{2}-2c^{2}(f^{\prime})^{2}=S(f)-2c^{2}(f^{\prime})^{2}.$
On the other hand, the two linearly independent local solutions found in Example 4.3 are $y_{1}=e^{cf}$ and $y_{2}=e^{-cf}$, hence
$w:=\frac{y_{1}}{y_{2}}=e^{2cf}.$
We now compute $S(w)$ in two different ways and compare it with the quantum Schwarzian.
Method 1: from the differential equation. By Proposition 2.1, the ratio $w=y_{1}/y_{2}$ satisfies
$S(w)=2\bigl(p^{\prime}-p^{2}-q\bigr)=\mathcal{S}_{p,q}.$
Method 2: by the Schwarzian chain rule. Using the Schwarzian chain rule $S(h\circ f)=(S(h)\circ f)(f^{\prime})^{2}+S(f)$ and the direct computation $S(e^{2ct})=-2c^{2}$, we obtain
$S(w)=S(e^{2cf})=S(f)-2c^{2}(f^{\prime})^{2}.$
This agrees with the expression for $\mathcal{S}_{p,q}$ above, and thus recovers the identity $S(y_{1}/y_{2})=\mathcal{S}_{p,q}$ in this concrete example.
Theorem 8.9.
Let AA be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. Suppose $g$ is an $\operatorname{End}(V)$-valued quantum 0-differential such that $\det(g_{i}(f))\not\equiv 0$ for every chart $U_{i}$ and every $f\in\mathcal{M}(U_{i})$, and assume that $g$ satisfies the second-order equation
$\nabla_{A}^{2}g=qg\qquad\bigl(\text{equivalently, }g^{\prime\prime}=2Ag^{\prime}+qg\bigr).$
Then for any two invertible solutions $g_{1},g_{2}$ of this equation, the connections $g_{1}^{-1}\bullet A$ and $g_{2}^{-1}\bullet A$ are classically equivalent.
Proof.
By Lemma 7.4, for any solution $g$ one has
$F_{g^{-1}\bullet A}=g^{-1}(F_{A}-q)g.$
If $g_{1}$ and $g_{2}$ are solutions, set $h:=g_{2}^{-1}g_{1}$ on the open subset where both are invertible. Then
$F_{g_{1}^{-1}\bullet A}=h^{-1}F_{g_{2}^{-1}\bullet A}h,$
which is precisely the condition of classical equivalence. ∎
Corollary 8.10.
Let AA be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. Let $g_{1}$ and $g_{2}$ be two invertible (classical) $\mathrm{GL}(V)$-valued solutions of
$g^{\prime\prime}=2Ag^{\prime}+qg.$
Then the connections $g_{1}^{-1}\bullet A$ and $g_{2}^{-1}\bullet A$ are equivalent in the sense of Definition 6.10. In particular, on each chart $U_{i}$ the coefficients of the characteristic polynomial of $F_{g^{-1}\bullet A,\,i}$ are independent of the choice of an invertible $\mathrm{GL}(V)$-valued solution $g$ of $g^{\prime\prime}=2Ag^{\prime}+qg$.
Moreover, since $F_{g^{-1}\bullet A}$ satisfies the same Schwarzian coordinate–change anomaly as any quantum curvature, these coefficients define invariants of the equation on $X$ in the same sense as the curvature (namely, their local expressions transform consistently on overlaps).
Proof.
Set $h:=g_{2}^{-1}g_{1}$ on the open subset where both solutions are pointwise invertible. By Lemma 8.9 one has
$F_{g_{1}^{-1}\bullet A}=h^{-1}F_{g_{2}^{-1}\bullet A}h.$
By Definition 6.10 this conjugacy of curvatures implies that the connections $g_{1}^{-1}\bullet A$ and $g_{2}^{-1}\bullet A$ are equivalent.
Finally, $F_{g_{1}^{-1}\bullet A}$ and $F_{g_{2}^{-1}\bullet A}$ are pointwise conjugate matrices, hence have the same characteristic polynomial. This proves that the coefficients of the characteristic polynomial of $F_{g^{-1}\bullet A}$ do not depend on the choice of $g$. ∎
Remark 8.11.
Let $g$ be an invertible (classical) $\mathrm{GL}(V)$-valued solution of the quantum second-order equation
$g^{\prime\prime}=2Ag^{\prime}+qg,$
and set
$B:=F_{g^{-1}\bullet A}.$
By Corollary 8.10, the conjugacy class of $B$ is independent of the choice of the solution $g$. Hence, for every integer $1\leq r\leq\dim(V)$, the functions $\operatorname{tr}(B^{r})$ (equivalently, the coefficients of the characteristic polynomial of $B$) are well-defined invariants of the differential equation, independent of $g$.
Moreover, Lemma 7.4 gives $B=g^{-1}(F_{A}-q)g$, and therefore for every $1\leq r\leq\dim(V)$ one has
$\operatorname{tr}(B^{r})=\operatorname{tr}\bigl((F_{A}-q)^{r}\bigr).$
In particular, these invariants are well defined on each chart and on $X$ (in the same sense as the curvature): under a change of coordinate the curvature acquires a Schwarzian anomaly, while $q$ transforms tensorially, so $F_{A}-q$ (and hence $B$) has the same Schwarzian anomaly as $F_{A}$.
Following Chern, we call the quantities $\operatorname{tr}(B^{r})$ (or, equivalently, the coefficients of the characteristic polynomial of $B$) the quantum characteristic invariants of the equation.
Part III Modularity
11. Modular forms and modular connections
Any hyperbolic Riemann surface, and in particular any Riemann surface of genus greater than one, can be realized as a quotient
$X=\Gamma\backslash\mathbb{H}$,
where $\mathbb{H}$ is the upper half-plane and $\Gamma\subset\operatorname{SL}_{2}(\mathbb{R})$ is a Fuchsian subgroup acting properly discontinuously.
Recall that an element
$\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{R})$
acts on $z\in\mathbb{H}$ by the Möbius transformation
$\gamma(z)=\frac{az+b}{cz+d}.$
Thus a meromorphic $m$-differential $f(z)(dz)^{m}$ on $X$ corresponds to a meromorphic function $f$ on $\mathbb{H}$ such that
$f(\gamma(z))\,(d\gamma(z))^{m}=f(z)\,dz^{m}\qquad(\gamma\in\Gamma).$
Since $d\gamma(z)=(cz+d)^{-2}dz$, we must have
(11.1) $f(\gamma z)=(cz+d)^{2m}f(z).$
Throughout this section, for a meromorphic function $h$ on $\mathbb{H}$ we write $h^{\prime}:=Dh:=\frac{1}{2\pi i}\frac{dh}{dz}$. For a Möbius transformation $\gamma(z)=\frac{az+b}{cz+d}$ we continue to write $\gamma^{\prime}(z)$ and $\gamma^{\prime\prime}(z)$ for the ordinary derivatives with respect to $z$ (so $\gamma^{\prime}(z)=(cz+d)^{-2}$).
Motivated by (11.1), for any integer $k$ and complex vector space $V$ we set
$M_{k}(\Gamma,V):=\left\{\psi:\mathbb{H}\to V\ \text{meromorphic}:\psi(\gamma z)=(cz+d)^{k}\psi(z)\ \text{for all }\gamma\in\Gamma\right\}.$
We refer to elements of $M_{k}(\Gamma,V)$ as $V$-valued meromorphic modular forms of weight $k$ for $\Gamma$.
In particular, weight 0 forms are precisely $\Gamma$-invariant meromorphic functions, hence
$M_{0}(\Gamma,\mathbb{C})\cong\mathcal{M}(X),$
and more generally $M_{0}(\Gamma,V)$ identifies with meromorphic $V$-valued functions on $X$.
Remark 11.1 (Classical $m$-differentials).
Let $V$ be finite-dimensional. A classical meromorphic $m$-differential on $X$ with values in $\operatorname{End}(V)$ may be represented on $\mathbb{H}$ as an $\operatorname{End}(V)$-valued meromorphic function $\phi$ such that $\phi(z)(dz)^{m}$ is $\Gamma$-invariant. Equivalently,
$\phi(\gamma z)=(cz+d)^{2m}\phi(z),\qquad\gamma\in\Gamma,$
so $\phi\in M_{2m}(\Gamma,\operatorname{End}(V))$. Conversely, every $\phi\in M_{2m}(\Gamma,\operatorname{End}(V))$ determines a classical $\operatorname{End}(V)$-valued $m$-differential on $X$.
11.1. Global quantum m-differentials
Definition 11.2.
[…]
$\Psi:M_{0}(\Gamma,V)\longrightarrow M_{2m}(\Gamma,V).$
This definition captures the basic idea that a quantum $m$-differential should send meromorphic functions on $X$ (weight 0) to meromorphic $m$-differentials on $X$ (weight $2m$).
To recover the local notion of quantum $m$-differential from Section 2 using an atlas whose transition maps lie in $\Gamma$, one may impose the following stronger equivariance condition.
Definition 11.3.
Let $m\in\mathbb{Z}$. For a linear operator
$\Psi:\mathcal{M}(\mathbb{H},V)\longrightarrow\mathcal{M}(\mathbb{H},V)$
and $\gamma\in\Gamma$, define the transformed operator
$(\Psi|_{m}\gamma)(f):=(\gamma^{\prime})^{m}\bigl(\Psi(f\circ\gamma^{-1})\bigr)\circ\gamma.$
We say that $\Psi$ is $\Gamma$-equivariant of weight $m$ if
$\Psi|_{m}\gamma=\Psi\qquad\text{for all }\gamma\in\Gamma.$
Proposition 11.4.
If $\Psi$ is $\Gamma$-equivariant of weight $m$, then it induces a global quantum $m$-differential in the sense of the first definition, i.e.
$\Psi\bigl(M_{0}(\Gamma,V)\bigr)\subseteq M_{2m}(\Gamma,V).$
Proof.
Let $f\in M_{0}(\Gamma,V)$, so $f\circ\gamma^{-1}=f$. Since $\Psi|_{m}\gamma=\Psi$, we have
$\Psi(f)=(\Psi|_{m}\gamma)(f)=(\gamma^{\prime})^{m}\,\Psi(f)\circ\gamma.$
Equivalently,
$\Psi(f)\circ\gamma=(\gamma^{\prime})^{-m}\Psi(f)=(cz+d)^{2m}\Psi(f),$
because $\gamma^{\prime}(z)=(cz+d)^{-2}$ for $\gamma\in\operatorname{SL}_{2}(\mathbb{R})$. Thus $\Psi(f)\in M_{2m}(\Gamma,V)$. ∎
Remark 11.5.
In the classical situation, multiplication by an $\operatorname{End}(V)$-valued $m$-differential provides an example of the above. Indeed, if $\phi\in M_{2m}(\Gamma,\operatorname{End}(V))$, then
$(\Psi_{\phi}f)(z):=\phi(z)f(z)$
defines a $\Gamma$-equivariant operator of weight $m$, hence a global quantum $m$-differential.
11.2. Quantum modular connections
Let $e\in\mathbb{C}$ and let $V$ be a finite-dimensional complex vector space. Recall that for
$\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{R})$
one has
$\gamma(z)=\frac{az+b}{cz+d},\qquad\gamma^{\prime}(z)=\frac{1}{(cz+d)^{2}}.$
Definition 11.6.
A quantum modular connection for $\Gamma$ with values in $V$ and eccentricity $e$ is a linear operator
$A:\mathcal{M}(\mathbb{H},V)\longrightarrow\mathcal{M}(\mathbb{H},V)$
such that for every $\gamma\in\Gamma$ one has
$A=A\|_{e}\gamma,$
where the transformed operator $A\|_{e}\gamma$ is defined by
$(A\|_{e}\gamma)(f):=\gamma^{\prime}\bigl(A(f\circ\gamma^{-1})\bigr)\circ\gamma+\frac{e}{2\pi i}\frac{\gamma^{\prime\prime}}{\gamma^{\prime}}f,\qquad f\in\mathcal{M}(\mathbb{H},V).$
Remark 11.7.
If AA is a quantum modular connection and $f\in M_{0}(\Gamma,V)$ is $\Gamma$-invariant, then $A(f)$ transforms by the inhomogeneous rule
$A(f)(\gamma z)=(cz+d)^{2}A(f)(z)+\frac{2e}{2\pi i}c(cz+d)f(z),\qquad\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma,$
[…]
11.3. Classical modular connections
Definition 11.8.
A classical modular connection for $\Gamma$ with values in $\operatorname{End}(V)$ and eccentricity $e$ is a meromorphic function
$a:\mathbb{H}\longrightarrow\operatorname{End}(V)$
such that for every $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma$,
$a(\gamma z)=(cz+d)^{2}a(z)+\frac{2e}{2\pi i}c(cz+d)I,$
where $I\in\operatorname{End}(V)$ denotes the identity endomorphism.
Remark 11.9.
If $a_{1}$ and $a_{2}$ are classical modular connections of the same eccentricity $e$, then their difference is a weight 22 matrix modular form:
$(a_{2}-a_{1})(\gamma z)=(cz+d)^{2}(a_{2}-a_{1})(z),$
so $a_{2}-a_{1}\in M_{2}(\Gamma,\operatorname{End}(V))$. Thus the space of classical modular connections of eccentricity $e$ is an affine space (a torsor) over $M_{2}(\Gamma,\operatorname{End}(V))$.
Proposition 11.10.
Let $a$ be a classical modular connection for $\Gamma$ with values in $\operatorname{End}(V)$ and eccentricity $e$. Define an operator
$A:\mathcal{M}(\mathbb{H},V)\longrightarrow\mathcal{M}(\mathbb{H},V),\qquad A(f):=a\cdot f.$
[…]
Proof.
Let $\gamma\in\Gamma$ and $f\in\mathcal{M}(\mathbb{H},V)$. Since $A(f\circ\gamma^{-1})=a(f\circ\gamma^{-1})$, we have
$\gamma^{\prime}\bigl(A(f\circ\gamma^{-1})\bigr)\circ\gamma=\gamma^{\prime}(a\circ\gamma)f.$
Hence
$(A\|_{e}\gamma)(f)=\bigl(\gamma^{\prime}(a\circ\gamma)+\frac{e}{2\pi i}\frac{\gamma^{\prime\prime}}{\gamma^{\prime}}I\bigr)f.$
Thus $A=A\|_{e}\gamma$ for all $\gamma$ is equivalent to
$a(z)=\gamma^{\prime}(z)a(\gamma z)+e\frac{\gamma^{\prime\prime}(z)}{\gamma^{\prime}(z)}I,$
which, after rewriting, is precisely the defining transformation rule for $a$. The $\mathcal{O}_{X}$-linearity is immediate. ∎
11.4. Modular covariant derivative
Let $e\neq 0$ and let $a:\mathbb{H}\to\operatorname{End}(V)$ be a classical modular connection for $\Gamma$ of eccentricity $e$.
Definition $11.$
For $\psi\in M_{k}(\Gamma,V)$ we define the modular covariant derivative associated to $a$ by
$\nabla_{a}(\psi):=\psi^{\prime}-\frac{k}{2e}a\psi,$
where $^{\prime}=D$ is the normalized derivation $D=\frac{1}{2\pi i}\frac{d}{dz}$ and $a$ acts on $V$ via the natural action of $\operatorname{End}(V)$.
Proposition 11.12.
The operator in the definition above is well-defined and satisfies
$\nabla_{a}:M_{k}(\Gamma,V)\longrightarrow M_{k+2}(\Gamma,V).$
Proof.
Let $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma$. Since $\psi\in M_{k}(\Gamma,V)$, we have
$\psi(\gamma z)=(cz+d)^{k}\psi(z).$
Differentiating and using $\gamma^{\prime}(z)=(cz+d)^{-2}$ yields
$\psi^{\prime}(\gamma z)=(cz+d)^{k+2}\psi^{\prime}(z)+\frac{k}{2\pi i}c(cz+d)^{k+1}\psi(z).$
On the other hand, the transformation rule for $a$ gives
$a(\gamma z)\psi(\gamma z)=(cz+d)^{k+2}a(z)\psi(z)+\frac{2e}{2\pi i}c(cz+d)^{k+1}\psi(z).$
Therefore
$(\nabla_{a}\psi)(\gamma z)=\psi^{\prime}(\gamma z)-\frac{k}{2e}a(\gamma z)\psi(\gamma z)=(cz+d)^{k+2}\left(\psi^{\prime}(z)-\frac{k}{2e}a(z)\psi(z)\right),$
since the terms proportional to $c(cz+d)^{k+1}\psi(z)$ cancel. Thus $\nabla_{a}\psi\in M_{k+2}(\Gamma,V)$. ∎
11.5. The Serre derivative
We now specialize to $\Gamma=\operatorname{SL}_{2}(\mathbb{Z})$. Using the normalized derivation
$D=\frac{1}{2\pi i}\frac{d}{dz},$
we write $\psi^{\prime}:=D\psi$ in this and the subsequent subsections (thus $d\psi/dz=2\pi i\,\psi^{\prime}$).
For $m\geq 2$, the Eisenstein series of weight $2m$ is
$E_{2m}(z)=1-\frac{4m}{B_{2m}}\sum_{n\geq 1}\sigma_{2m-1}(n)q^{n},\qquad q=e^{2\pi iz},$
where $B_{2m}$ is the Bernoulli number and $\sigma_{r}(n)=\sum_{d\mid n}d^{r}$. For $m=1$ the above series is no longer absolutely convergent; the holomorphic function defined by the Fourier expansion
$E_{2}(z)=1-24\sum_{n\geq 1}\sigma_{1}(n)q^{n}$
is the (holomorphic) Eisenstein series of weight 22. Unlike $E_{2m}$ for $m\geq 2$, the function $E_{2}$ is not a modular form. Instead it satisfies the quasi-modular transformation law
$E_{2}(\gamma z)=(cz+d)^{2}E_{2}(z)+\frac{12}{2\pi i}c(cz+d),\qquad\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z}).$
More generally, a holomorphic function $f:\mathbb{H}\to\mathbb{C}$ is called a quasi-modular form of weight $k$ and depth $\leq p$ (for $\operatorname{SL}_{2}(\mathbb{Z})$) if for every $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z})$ one has
$f(\gamma z)=(cz+d)^{k}\sum_{r=0}^{p}f_{r}(z)\Bigl(\frac{c}{cz+d}\Bigr)^{r}$
for some holomorphic functions $f_{r}$ on $\mathbb{H}$. The graded ring of quasi-modular forms for $\operatorname{SL}_{2}(\mathbb{Z})$ is generated by $E_{2},E_{4},E_{6}$; see [21].
The Ramanujan discriminant modular form of weight $12$ is
$\Delta(z)=q\prod_{n\geq 1}(1-q^{n})^{24},\qquad\Delta=\frac{E_{4}^{3}-E_{6}^{2}}{1728}.$
From the transformation law
$\Delta(\gamma z)=(cz+d)^{12}\Delta(z),$
one computes
$\frac{\Delta^{\prime}}{\Delta}(\gamma z)=(cz+d)^{2}\frac{\Delta^{\prime}}{\Delta}(z)+\frac{12}{2\pi i}c(cz+d),$
where here denotes $D$. Hence the function
$a(z)=\frac{1}{6}\frac{\Delta^{\prime}(z)}{\Delta(z)}I$
satisfies
$a(\gamma z)=(cz+d)^{2}a(z)+\frac{2}{2\pi i}c(cz+d)I,$
so $a$ is a classical modular connection of eccentricity $e=1$. By the modular covariant derivative definition, the associated covariant derivative on $M_{k}(\Gamma,V)$ is
$\nabla_{a}(\psi)=\psi^{\prime}-\frac{k}{2}a\psi.$
Substituting $a=\frac{1}{6}\frac{\Delta^{\prime}}{\Delta}=\frac{1}{6}E_{2}I$ yields
$\nabla_{a}(\psi)=\psi^{\prime}-\frac{k}{12}E_{2}\psi,$
which is precisely the Serre derivative introduced in [19].
11.6. Ramanujan identities
For $\Gamma=\operatorname{SL}_{2}(\mathbb{Z})$, the Eisenstein series $E_{2},E_{4},E_{6}$ satisfy the Ramanujan differential identities (see Zagier [21]):
$E^{\prime}_{2}=\frac{E_{2}^{2}-E_{4}}{12},\qquad E^{\prime}_{4}=\frac{E_{2}E_{4}-E_{6}}{3},\qquad E^{\prime}_{6}=\frac{E_{2}E_{6}-E_{4}^{2}}{2}.$
Remark 11.13.
[…]
$F_{A}=A^{\prime}-\frac{1}{2e}A^{2}$
defines a well-defined modular 22-differential on $X$. Indeed, the anomaly term involves the Schwarzian $S(\lambda)$ of the transition function, and $S(\lambda)=0$ for Möbius changes of coordinates. If AA is a classical modular connection with values in $\operatorname{End}(V)$, then $F_{A}\in M_{4}(\Gamma,\operatorname{End}(V))$.
11.7. Modular Curvature
Example 11.14.
Let $f:\mathbb{H}\to V$ be an element of $M_{2m}(\Gamma,V)$, where $V$ is an $n$-dimensional complex vector space. Define the Wronskian
$\operatorname{Wr}(f)(z)=f(z)\wedge f^{\prime}(z)\wedge\cdots\wedge f^{(n-1)}(z)\in\bigwedge^{n}V,$
where derivatives are taken with respect to the normalized derivation $D$. As in Section 2, $\operatorname{Wr}(f)$ transforms as a meromorphic modular form of weight
N:=2mn+n(n1),N:=2mn+n(n-1),
so $\operatorname{Wr}(f)\in M_{N}(\Gamma,\bigwedge^{n}V)$. Since $\bigwedge^{n}V$ is one-dimensional, we may regard $\operatorname{Wr}(f)$ as a scalar modular form up to a choice of basis. A change of basis multiplies $\operatorname{Wr}(f)$ by $\det(T)$, hence the logarithmic derivative $\operatorname{Wr}(f)^{\prime}/\operatorname{Wr}(f)$ is independent of the choice of basis.
Define the scalar Maurer–Cartan connection
$A=\frac{2}{N}\frac{\operatorname{Wr}(f)^{\prime}}{\operatorname{Wr}(f)}.$
Then AA is a classical modular connection of eccentricity $e=1$. Consequently its curvature
$F_{A}=A^{\prime}-\frac{1}{2}A^{2}$
is a meromorphic modular form of weight 44. Thus we obtain a natural nonlinear operator
$M_{2m}(\Gamma,V)\longrightarrow M_{4}(\Gamma,\mathbb{C}),\qquad f\longmapsto F_{A}.$
11.8. The Chazy equation
In the classical Cartan theory of Lie-algebra-valued connections, the curvature $F=dA+\frac{1}{2}[A,A]$ satisfies the Bianchi identity $d_{A}F=0$. In our one-dimensional framework, the curvature of a connection is the quadratic expression
$F_{A}=A^{\prime}-\frac{1}{2e}A^{2},$
which should be viewed as a curvature-type 22-differential. However, as we saw in Section 3, $F_{A}$ is not globally tensorial on an arbitrary Riemann surface: under changes of coordinate it acquires an anomaly term proportional to the Schwarzian derivative of the transition map. Thus one should not expect a literal analogue of the classical Bianchi identity in general. Chazy introduced in [1] a class of nonlinear third-order differential equations whose general solutions have fixed critical points. A distinguished example is the Chazy equation
(11.2) $2u^{\prime\prime\prime}-2uu^{\prime\prime}+3(u^{\prime})^{2}=0,$
where, in this subsection, primes denote the normalized derivation $D$. It is a classical fact (see Zagier [21]) that the Eisenstein series $E_{2}$ satisfies (11.2). We now derive this from the modular connection $A=\frac{1}{6}\frac{\Delta^{\prime}}{\Delta}$.
Lemma 11.15 ( Chazy equation).
Let $\Gamma=\operatorname{SL}_{2}(\mathbb{Z})$ and write $f^{\prime}:=Df$ for $D=\frac{1}{2\pi i}\frac{d}{dz}$. Let
$A=\frac{1}{6}\frac{\Delta^{\prime}}{\Delta}=\frac{1}{6}E_{2},$
so that AA is a scalar classical modular connection of eccentricity $e=1$. Let $F_{A}=A^{\prime}-\frac{1}{2}A^{2}$ be the curvature. Then AA satisfies the Chazy equation
$\nabla_{A}^{2}F_{A}+12F_{A}^{2}=0,$
Part IV Periods
In this section we apply the ideas developed so far to study periods of holomorphic 11-forms on a one-parameter family of compact Riemann surfaces of fixed genus $g$, generalizing the formula of Dedekind mentioned in the introduction. Differential equations satisfied by periods of holomorphic 11-forms often fall under the title of Picard–Fuchs equations: these are scalar linear ODEs of order $2g$ obtained from the Gauss–Manin system. While extremely useful, such scalar equations are not well suited for the Schwarzian point of view pursued in this paper. For $g>1$ we therefore propose a different approach which produces a canonical second order ODE with $g\times g$ matrix coefficients in $\operatorname{End}(E)$, and from there we obtain new insight about periods using the matrix Schwarzian derivative. Such a one-parameter family, roughly speaking, either corresponds to a non-constant holomorphic map from a Riemann surface $S$ to the moduli space of curves of genus $g$, $\mathcal{M}_{g}$, or equivalently to a holomorphic fibration $\pi\colon X\to S$ from a complex surface $X$ to $S$ whose fibers $X_{s}=\pi^{-1}(s)$ are Riemann surfaces of genus $g$.
For a compact Riemann surface $C$ of genus $g$, its first cohomology group $H^{1}(C,\mathbb{C})$ carries a natural Hodge decomposition
$H^{1}(C,\mathbb{C})\;=\;H^{1,0}(C)\oplus H^{0,1}(C),$
where $H^{1,0}(C)\simeq H^{0}(C,K_{C})$ is the space of holomorphic 11-forms and $H^{0,1}(C)\simeq\overline{H^{1,0}(C)}$ is the space of anti-holomorphic 11-forms.
Associated to a “generic” family $\pi\colon X\to S$ (to be made precise later), or equivalently to the induced moduli map
$\phi\colon S\to\mathcal{M}_{g},\qquad s\longmapsto[X_{s}],$
we will construct a classical connection
$A=A_{\phi}$
of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(E)$, where
$E:=\pi_{*}\Omega^{1}_{X/S}$
is the Hodge bundle whose fiber at $s\in S$ is $H^{1,0}(X_{s})$. We will also construct a classical 22-differential
$q=q_{\phi}$
with values in $\operatorname{End}(E)$. The associated second order ODE with matrix coefficients
$\Psi^{\prime\prime}=2A\Psi^{\prime}+q\Psi$
will then encode the variation of periods: for a locally constant homology class $\gamma\in H_{1}(X_{s},\mathbb{Z})$ and a holomorphic 11-form $\omega_{s}\in H^{1,0}(X_{s})$ varying holomorphically with $s$, the period function
$s\longmapsto\int_{\gamma_{s}}\omega_{s}$
arises as a component of a solution $\Psi$ of the above equation.
We briefly recall that the traditional Picard–Fuchs approach starts from the Gauss–Manin system (a first order linear system of rank $2g$) and produces a scalar linear ODE of order $2g$ by choosing a cyclic vector in the sense of Katz. However, for $g>1$ there is in general no canonical choice of such a cyclic vector (or of a distinguished scalar period), so the resulting scalar equation depends on auxiliary choices. By contrast, the second order ODE with coefficients in $\operatorname{End}(E)$ that we construct below is canonical: it is determined intrinsically by the family (under an explicit genericity hypothesis) and is compatible with the gauge–theoretic formalism developed in the previous sections.
11.9. Dedekind’s formula
We illustrate the preceding discussion in the case of elliptic curves ($g=1$), recovering Dedekind’s Schwarzian differential equation from a Picard–Fuchs computation. Throughout this subsection, $j$ denotes the normalized Hauptmodul on $X(1)\simeq\mathbb{P}^{1}$ used in the introduction, so that the three special points correspond to $j=0,1,\infty$.
The universal family. Consider the one-parameter family of smooth plane cubic curves
(11.3) $E_{j}:\qquad y^{2}=4x^{3}-g(j)\,x-g(j),\qquad g(j)=\frac{27j}{j-1},$
over $U:=\mathbb{P}^{1}\setminus\{0,1,\infty\}$. For the Weierstrass equation $y^{2}=4x^{3}-g_{2}x-g_{3}$, one has discriminant $\Delta=g_{2}^{3}-27g_{3}^{2}$ and normalized modular invariant
$j=\frac{g_{2}^{3}}{\Delta}.$
In our family $g_{2}=g_{3}=g(j)$, hence
$\Delta=g(j)^{2}\bigl(g(j)-27\bigr),\qquad\frac{g(j)^{3}}{\Delta}=\frac{g(j)}{g(j)-27}=j,$
so (11.3) indeed gives the universal elliptic curve over the $j$-line in this normalization (cf. [5]).
Gauss–Manin system. Let
$\omega=\frac{dx}{y},\qquad\eta=\frac{x\,dx}{y}$
be the standard holomorphic differential and a differential of the second kind on $E_{j}$. Then $\{\omega,\eta\}$ is a basis of $H^{1}_{\mathrm{dR}}(E_{j})$. Differentiating in the parameter $g$ (and using $y^{2}=4x^{3}-gx-g$) gives
$\partial_{g}\omega=\frac{x+1}{2}\frac{dx}{y^{3}},\qquad\partial_{g}\eta=\frac{x(x+1)}{2}\frac{dx}{y^{3}}.$
By a standard reduction step (subtracting exact differentials of the form $d(H/y)$; see [15]), one can express these derivatives in the basis $\{\omega,\eta\}$.
Lemma 11.16.
In $H^{1}_{\mathrm{dR}}(E_{j})$ one has
(11.4) $\frac{d}{dg}\binom{\omega}{\eta}=\begin{pmatrix}\displaystyle\frac{18-g}{4g(g-27)}&\displaystyle-\frac{3}{2g(g-27)}\\[8.0pt] \displaystyle\frac{1}{8(g-27)}&\displaystyle\frac{g-18}{4g(g-27)}\end{pmatrix}\binom{\omega}{\eta}$,
where $g=g(j)$ as in (11.3).
Proof.
We sketch the reduction for $\partial_{g}\omega$; the computation for $\partial_{g}\eta$ is similar. Using $d(H/y)=H^{\prime}\,dx/y-\frac{1}{2}H\,(12x^{2}-g)\,dx/y^{3}$ and the identity $dx/y=(4x^{3}-gx-g)\,dx/y^{3}$, a direct calculation shows that
$\partial_{g}\omega=\frac{18-g}{4g(g-27)}\,\omega-\frac{3}{2g(g-27)}\,\eta+d\!\left(\frac{H}{y}\right),$
where one may take
$H(x)=\frac{1}{g-27}\Bigl(\frac{3}{g}x^{2}+\frac{18-g}{2g}x-\frac{1}{2}\Bigr).$
Similarly, one checks that
$\partial_{g}\eta=\frac{1}{8(g-27)}\,\omega+\frac{g-18}{4g(g-27)}\,\eta+d\!\left(\frac{H_{2}}{y}\right),$
where
$H_{2}(x)=\frac{1}{g-27}\Bigl(\frac{18-g}{2g}x^{2}+\frac{1}{4}x+\frac{3}{4}\Bigr).$
which yields the second row of (11.4).
Since exact forms integrate to zero over closed cycles, this yields the first row of (11.4). The second row follows from an analogous reduction of $\partial_{g}\eta$. ∎
Let $\gamma$ be a locally constant cycle in $H_{1}(E_{j},\mathbb{Z})$. Integrating (11.4) over $\gamma$ yields a first order linear system for the period vector
$\Pi_{\gamma}(g)=\binom{\int_{\gamma}\omega}{\int_{\gamma}\eta}$.
Eliminating the second component gives a scalar second-order equation for $y(g):=\int_{\gamma}\omega$.
Proposition 11.17 (Picard–Fuchs equation in the jj–coordinate).
The periods $y(j)=\int_{\gamma}\omega$ satisfy the Picard–Fuchs equation
(11.5) $\frac{d^{2}y}{dj^{2}}+\frac{1}{j}\frac{dy}{dj}+\frac{31j-4}{144j^{2}(1-j)^{2}}\,y=0.$
Proof.
Let $\gamma$ be a locally constant cycle and set
$y(g):=\int_{\gamma}\omega,\qquad z(g):=\int_{\gamma}\eta.$
Integrating (11.4) gives a first-order system
$\frac{d}{dg}\binom{y}{z}=\begin{pmatrix}a&b\\ c&-a\end{pmatrix}\binom{y}{z},\qquad\begin{aligned} a&=\frac{18-g}{4g(g-27)},\\ b&=-\frac{3}{2g(g-27)},\\ c&=\frac{1}{8(g-27)}.\end{aligned}$
From the first equation we have $z=(y^{\prime}-ay)/b$. Differentiating this identity and using the second equation to eliminate $z^{\prime}$ yields the scalar equation
$\frac{d^{2}y}{dg^{2}}+\frac{2g-27}{g(g-27)}\frac{dy}{dg}+\frac{3(g+4)}{16g^{2}(g-27)}\,y=0.$
Finally, using the change of variable $g=\frac{27j}{j-1}$ and the chain rule converts this into (11.5). This is the equation stated in the introduction. ∎
Dedekind’s Schwarzian equation. Let $\omega_{1}(j),\omega_{2}(j)$ be two linearly independent solutions of (11.5) and set $\tau(j)=\omega_{2}(j)/\omega_{1}(j)$. Changing the ordered basis $(\omega_{1},\omega_{2})$ acts on $\tau$ by a Möbius transformation, hence $S(\tau)$ is well defined. For a choice of period basis the ambiguity is the natural action of $\operatorname{SL}_{2}(\mathbb{Z})$ on $\tau$. In particular, $\tau(j)$ depends only on the period lattice and gives a uniformizing parameter on $\mathbb{H}$. Putting (11.5) into normal form by removing the first derivative term, one finds that $\tau$ satisfies a Schwarzian differential equation. More precisely, from (11.5) one computes
$S(\tau)(j)=2\left(b-\frac{1}{2}a^{\prime}-\frac{1}{4}a^{2}\right),\qquad a=\frac{1}{j},\quad b=\frac{31j-4}{144j^{2}(1-j)^{2}},$
and hence
(11.6) $S(\tau)(j)=\frac{3}{8(1-j)^{2}}+\frac{4}{9j^{2}}+\frac{23}{72j(1-j)}.$
This is precisely the formula found by Dedekind [5]; compare the discussion in the introduction. In the language of Section 4, rewriting (11.5) in the form $y^{\prime\prime}=2Ay^{\prime}+qy$ with
$A=-\frac{1}{2j},\qquad q=-\frac{31j-4}{144j^{2}(1-j)^{2}},$
one obtains $2(F_{A}-q)=S(\tau)(j)$. We now proceed to extend this construction to higher genus $g>1$, where the ODE remains of second order but the coefficients become $g\times g$ matrices.
12. Higher genus periods
On the moduli space $\mathcal{M}_{g}$ there is a canonical holomorphic vector bundle $\mathbb{E}$ of rank $g$ (the Hodge bundle) whose fiber at a point $C]\in\mathcal{M}_{g}$ is
$\mathbb{E}_{[C]}\;\cong\;H^{1,0}(C)\;\cong\;H^{0}(C,K_{C}).$
In algebro–geometric terms, if $\pi\colon\mathcal{C}_{g}\to\mathcal{M}_{g}$ denotes the universal curve (viewed as a stack), then
$\mathbb{E}=\pi_{*}\omega_{\pi},$
where $\omega_{\pi}$ is the relative dualizing sheaf [12]. For a family $\pi\colon X\to S$ with induced moduli map $\phi\colon S\to\mathcal{M}_{g}$, the bundle $E=\pi_{*}\Omega^{1}_{X/S}$ introduced above is naturally identified with the pullback $\phi^{*}\mathbb{E}$.
Alongside $\mathbb{E}$ one has the rank $2g$ local system
$\mathbb{H}:=R^{1}\pi_{*}\mathbb{C}$
on $S$, whose fiber at $s\in S$ is $H^{1}(X_{s},\mathbb{C})$. The associated holomorphic vector bundle
$\mathcal{H}:=\mathbb{H}\otimes_{\mathbb{C}}\mathcal{O}_{S}\cong H^{1}_{\mathrm{dR}}(X/S)$
admits the Gauss–Manin connection
$\nabla^{\mathrm{GM}}\colon\mathcal{H}\longrightarrow\mathcal{H}\otimes\Omega^{1}_{S},\qquad(\nabla^{\mathrm{GM}})^{2}=0,$
which is a canonical integrable (flat) connection encoding the parallel transport of cohomology classes in families [15]. The Hodge decomposition of each fiber $H^{1}(X_{s},\mathbb{C})\cong H^{1,0}(X_{s})\oplus H^{0,1}(X_{s})$ is not preserved by $\nabla^{\mathrm{GM}}$; instead one has Griffiths transversality $\nabla^{\mathrm{GM}}(E)\subset(\mathcal{H}/E)\otimes\Omega^{1}_{S}$, which reflects the variation of complex structure.
Fix a simply connected open set $U\subset S$ with coordinate $t$. Choosing a flat basis of cycles $\gamma_{1},\dots,\gamma_{2g}$ for the local system $R_{1}\pi_{*}\mathbb{Z}$ and a holomorphic frame $\omega_{1},\dots,\omega_{g}$ of $E|_{U}$, the period matrix
$\Pi(t):=\bigl(\,\int_{\gamma_{j}(t)}\omega_{i}(t)\,\bigr)_{1\leq i\leq g,\,1\leq j\leq 2g}$
satisfies a first order linear differential system (the Gauss–Manin system) obtained by expressing $\nabla^{\mathrm{GM}}\omega_{i}$ in the chosen basis. Equivalently, the vector of all period functions is a horizontal section of $(\mathcal{H},\nabla^{\mathrm{GM}})$.
In the one-parameter case one may also extract from the rank $2g$ Gauss–Manin system a scalar linear ordinary differential equation of order $2g$ (a Picard–Fuchs equation) annihilating a given period function. Katz gives an algebraic construction of the Gauss–Manin connection and a method to pass from the first-order system to such a scalar equation by choosing a cyclic vector [14, 18]. This reduction is highly useful, but it is not canonical in general: the resulting scalar equation depends on the choice of cyclic vector (equivalently, on the elimination procedure). In this section we instead emphasize the canonical second order ODE with coefficients in $\operatorname{End}(E)$ attached to a generic family.
Definition 12.1.
Let $\pi\colon X\to S$ be a holomorphic family of compact Riemann surfaces of genus $g\geq 1$, and let $E=\pi_{*}\Omega^{1}_{X/S}$ be the Hodge bundle on $S$. On a simply connected open set $U\subset S$ with local coordinate $s$, let
$\theta_{\partial_{s}}\colon E|_{U}\longrightarrow(R^{1}\pi_{*}\mathcal{O}_{X})|_{U}$
denote the Higgs field (equivalently, the $(0,1)$-part of the Gauss–Manin connection along $\partial_{s}$). We say that the family $\pi$ is generic on $U$ if $\theta_{\partial_{s}}$ is an isomorphism on a dense open subset of $U$ (equivalently, if $\det(\theta_{\partial_{s}})$ is not identically zero on $U$). We say that $\pi$ is generic if it is generic on every sufficiently small coordinate chart $U\subset S$.
Remark 12.2.
Since $\operatorname{rank}(E)=\operatorname{rank}(R^{1}\pi_{*}\mathcal{O}_{X})=g$, the condition in Definition 12.1 is equivalent to requiring that for a generic point $s\in U$ one can choose a holomorphic frame $\omega_{1},\dots,\omega_{g}$ of $E|_{U}$ such that the classes of $\partial_{s}\omega_{1},\dots,\partial_{s}\omega_{g}$ span $H^{0,1}(X_{s})$. Equivalently, for a generic $s\in U$ the induced bilinear form
$(\omega,\eta)\longmapsto\langle\kappa_{s}(\partial_{s}),\,\omega\eta\rangle\qquad\bigl(\omega,\eta\in H^{1,0}(X_{s})\bigr)$
is nondegenerate, where $\kappa_{s}(\partial_{s})\in H^{1}(X_{s},T_{X_{s}})$ is the Kodaira–Spencer class.
Definition 12.3.
Let $\pi\colon X\to S$ be a generic family of compact Riemann surfaces of genus $g$ over a Riemann surface $S$, and let
$E:=\pi_{*}\Omega^{1}_{X/S}$
be the Hodge bundle on $S$. Let $\mathcal{H}=H^{1}_{\mathrm{dR}}(X/S)$ denote the relative de Rham bundle equipped with the Gauss–Manin connection $\nabla^{\mathrm{GM}}$. Fix a simply connected coordinate chart $U\subset S$ with local coordinate $s$ on which the genericity condition holds. Choosing a flat trivialization of $(\mathcal{H},\nabla^{\mathrm{GM}})$ over $U$, we identify covariant differentiation $\nabla^{\mathrm{GM}}_{\partial_{s}}$ with ordinary differentiation and write primes for $d/ds$.
Assume that $E|_{U}$ admits a holomorphic frame $\omega_{1},\dots,\omega_{g}$ such that the $2g$ sections
$\omega_{1},\dots,\omega_{g},\ \omega_{1}^{\prime}:=\partial_{s}\omega_{1},\dots,\partial_{s}\omega_{g}=:\omega_{g}^{\prime}$
form a frame of $\mathcal{H}|_{U}$ (equivalently, the matrix of the Higgs field $\theta_{\partial_{s}}$ in the frame $\omega_{1},\dots,\omega_{g}$ is pointwise invertible on a dense open subset of $U$). Let
$\omega:=\begin{pmatrix}\omega_{1}\\ \vdots\\ \omega_{g}\end{pmatrix}$.
Since $(\omega,\omega^{\prime})$ is a frame of $\mathcal{H}|_{U}$, there exist unique meromorphic matrix-valued functions
$A_{U},\ q_{U}\in\mathcal{M}(U,\operatorname{End}(\mathbb{C}^{g}))$
such that
[…]
We call $A_{U}$ and $q_{U}$ the local coefficient matrices attached to the family $\pi$ (or equivalently to the moduli map $\phi$) on $U$.
Lemma 12.4.
In the setting of Definition 12.3, the collections $\{A_{U}\}$ and $\{q_{U}\}$ glue to define a classical connection $A=A_{\phi}$ on $E$ of eccentricity $e=\tfrac{1}{2}$ and a classical 22-differential $q=q_{\phi}$ with values in $\operatorname{End}(E)$.
More precisely, let $(U_{i},s_{i},\omega^{(i)})$ and $(U_{j},s_{j},\omega^{(j)})$ be two choices as in Definition 12.3, and write $\lambda_{ij}=s_{i}\circ s_{j}^{-1}$ for the change of coordinate on $U_{i}\cap U_{j}$. Assume that the corresponding frames are related by a constant matrix $g_{ij}\in\operatorname{GL}_{g}(\mathbb{C})$,
$\omega^{(j)}=g_{ji}\,(\omega^{(i)}\circ\lambda_{ij}),\qquad g_{ji}=g_{ij}^{-1}.$
Then on $U_{i}\cap U_{j}$ one has
$A_{j}=g_{ji}\Bigl(\lambda^{\prime}_{ij}(A_{i}\circ\lambda_{ij})+\frac{1}{2}\frac{\lambda^{\prime\prime}_{ij}}{\lambda^{\prime}_{ij}}I\Bigr)g_{ij},\qquad q_{j}=(\lambda^{\prime}_{ij})^{2}\,g_{ji}\,(q_{i}\circ\lambda_{ij})\,g_{ij}.$
In particular, AA transforms as a (classical) connection of eccentricity $e$ and $q$ transforms as an $\operatorname{End}(E)$-valued 22-differential.
Proof.
Fix a point of $U_{i}\cap U_{j}$ and write $\lambda=\lambda_{ij}$. Set $p_{i}:=2A_{i}$. If we represent the same section in the $(U_{j},s_{j})$-chart by
$\widetilde{\omega}:=\omega^{(i)}\circ\lambda,$
then $\widetilde{\omega}$ satisfies a second-order equation of the form
$\widetilde{\omega}^{\prime\prime}=\widetilde{p}_{j}\,\widetilde{\omega}^{\prime}+\widetilde{q}_{j}\,\widetilde{\omega}$
with respect to the coordinate $s_{j}$. By Lemma 7.1 (applied componentwise to the vector–valued function $\widetilde{\omega}$), the coefficients are
$\widetilde{p}_{j}=\lambda^{\prime}(p_{i}\circ\lambda)+\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}I,\qquad\widetilde{q}_{j}=(\lambda^{\prime})^{2}(q_{i}\circ\lambda).$
Writing $\widetilde{p}_{j}=2\widetilde{A}_{j}$, we obtain
$\widetilde{A}_{j}=\lambda^{\prime}(A_{i}\circ\lambda)+\frac{1}{2}\frac{\lambda^{\prime\prime}}{\lambda^{\prime}}I.$
Finally, the chosen frame on $U_{j}$ is
$\omega^{(j)}=g_{ji}\,\widetilde{\omega},\qquad g_{ji}=g_{ij}^{-1},$
with $g_{ji}$ constant. Hence differentiation commutes with multiplication by $g_{ji}$, and $\omega^{(j)}$ satisfies
$(\omega^{(j)})^{\prime\prime}=2A_{j}(\omega^{(j)})^{\prime}+q_{j}\omega^{(j)}$
with
$A_{j}=g_{ji}\,\widetilde{A}_{j}\,g_{ij},\qquad q_{j}=g_{ji}\,\widetilde{q}_{j}\,g_{ij}.$
Substituting the expressions for $\widetilde{A}_{j}$ and $\widetilde{q}_{j}$ yields the stated transformation laws. ∎
13. A genus two family
We conclude the section with a concrete genus two example. Recall that every smooth curve of genus 22 is hyperelliptic, so hyperelliptic families provide a natural and sufficiently general testing ground for the matrix Schwarzian formalism.
Let $U\subset\mathbb{C}$ be a simply connected open set avoiding the discriminant locus, and consider the one–parameter family of genus 22 curves
(13.1) $X_{t}:\qquad y^{2}=x^{5}+t\,x^{2}+1,\qquad t\in U.$
The fiber $X_{t}$ is smooth if and only if
$D(t):=\operatorname{Res}_{x}(x^{5}+t\,x^{2}+1,\ 5x^{4}+2tx)=108t^{5}+3125$
is nonzero.
On $X_{t}$ the holomorphic 11-forms are spanned by
$\omega_{1}=\frac{dx}{y},\qquad\omega_{2}=\frac{x\,dx}{y}.$
Writing $\omega=(\omega_{1},\omega_{2})^{t}$, differentiation with respect to $t$ gives meromorphic differentials of the second kind
$\omega_{1}^{\prime}=-\frac{1}{2}\,\frac{x^{2}\,dx}{y^{3}},\qquad\omega_{2}^{\prime}=-\frac{1}{2}\,\frac{x^{3}\,dx}{y^{3}}.$
A direct computation in $H^{1}_{\mathrm{dR}}(X_{t})$ (for instance via Griffiths reduction) shows that the 44 classes
$\omega_{1},\ \omega_{2},\ \omega_{1}^{\prime},\ \omega_{2}^{\prime}$
form a basis for $H^{1}_{\mathrm{dR}}(X_{t})$ for $D(t)\neq 0$, so the family is generic in the sense of Definition 12.1 on $U$. Therefore, by Definition 12.3 there exist unique $2\times 2$ matrix–valued functions $A(t)$ and $q(t)$ such that
(13.2) $\omega^{\prime\prime}=2A(t)\,\omega^{\prime}+q(t)\,\omega.$
For the family (13.1) one finds explicitly
(13.3) $A(t)=\frac{1}{D(t)}\begin{pmatrix}-108t^{4}&-675t^{2}\\ 750t&-162t^{4}\end{pmatrix},\qquad q(t)=\frac{1}{D(t)}\begin{pmatrix}-27t^{3}&-\frac{275}{2}t\\[2.0pt] \frac{375}{2}&-33t^{3}\end{pmatrix}$.
Since $e=\tfrac{1}{2}$, the curvature of AA is $F_{A}=A^{\prime}-A^{2}$ and the (matrix) quantum Schwarzian associated to (13.2) is
$\mathcal{S}_{A,q}:=2(F_{A}-q).$
Using (13.3) one obtains the following $\operatorname{End}(\mathbb{C}^{2})$-valued rationalrational 22-differential:
(13.4) $\mathcal{S}_{A,q}(t)=\frac{1}{D(t)^{2}}\begin{pmatrix}t^{3}(5832t^{5}-1518750)&25t(4104t^{5}-303125)\\[3.0pt] 1125(3125-252t^{5})&t^{3}(-10368t^{5}-2831250)\end{pmatrix}$.
Remark 13.1.
The matrix Schwarzian $\mathcal{S}_{A,q}$ produces scalar invariants by taking characteristic combinations. In the present genus two case one obtains in particular the 22-differentials
$\operatorname{tr}(\mathcal{S}_{A,q})\qquad\text{and}\qquad\operatorname{tr}(\mathcal{S}_{A,q}^{2}),$
which are well defined on each chart and hence on $X$ in the same sense as the quantum curvature (they inherit the same Schwarzian anomaly under coordinate change). By Lemma 7.4 and Remark 8.11, these are the first instances of the quantum characteristic invariants attached to the second-order equation (13.2).
In genus one, Dedekind’s computation expresses the Schwarzian derivative of the ratio of two periods as $2(F_{A}-q)$ for a scalar equation; see (11.6). Lemma 7.4 shows that $2(F_{A}-q)$ is the intrinsic Schwarzian object associated to a second order equation, and the present construction may be viewed as its extension from elliptic curves to families of curves of arbitrary genus $g$, where the ODE remains of second order but the coefficients become $g\times g$ matrices and the resulting matrix Schwarzian yields canonical characteristic invariants. Finally, since every genus two curve is hyperelliptic, the family (13.1) is a natural test case for the higher genus theory.
14. Cubic threefolds in $\mathbb{P}^{4}$
In this section we illustrate how the second order non-abelian formalism developed in Section 7 and Part IV extends to a basic higher-dimensional moduli problem, namely smooth cubic threefolds in $\mathbb{P}^{4}$. While the Gauss–Manin connection for a one-parameter family of cubic threefolds is a first-order system of rank $10$ on $H^{3}$, our approach produces a canonical second order system on the rank-55 Hodge bundle $E=H^{2,1}$ with coefficients in $\operatorname{End}(E)$. For a natural deformation of the Fermat cubic, a large symmetry group forces this system to split into one-dimensional character spaces; in a symmetry-adapted basis the coefficient matrices become diagonal. We explain the geometric input (Hodge theory, residues, and the Gauss–Manin system), state the resulting diagonal second order ODE, and omit only the standard reduction computations.
14.1. Cubic threefolds
A cubic threefold is a smooth hypersurface $X\subset\mathbb{P}^{4}$ defined by a homogeneous polynomial $F(x_{0},\dots,x_{4})=0$ of degree 33. Two classical examples are the Fermat cubic threefold
$X_{\mathrm{F}}:\quad x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+x_{4}^{3}=0,$
and the Klein cubic threefold
$X_{\mathrm{K}}:\quad x_{0}^{2}x_{1}+x_{1}^{2}x_{2}+x_{2}^{2}x_{3}+x_{3}^{2}x_{4}+x_{4}^{2}x_{0}=0.$
Both are smooth and admit large automorphism groups.
The Hodge numbers of a smooth cubic threefold are classical:
$h^{3,0}(X)=h^{0,3}(X)=0,\qquad h^{2,1}(X)=h^{1,2}(X)=5,\qquad h^{1,1}(X)=1,$
so that
$H^{3}(X,\mathbb{C})=H^{2,1}(X)\oplus H^{1,2}(X),\qquad\dim H^{2,1}(X)=5.$
See for instance [3, 20].
Since $h^{3,0}=0$, the Hodge structure on $H^{3}(X)$ is of type $(2,1)+(1,2)$. Writing $M:=X$, the associated intermediate Jacobian can be presented as the complex torus
$J(M):=H^{1,2}(M)\big/H^{3}(M,\mathbb{Z}),$
where we view $H^{3}(M,\mathbb{Z})$ as a lattice in $H^{1,2}(M)$ via the projection $H^{3}(M,\mathbb{C})\twoheadrightarrow H^{1,2}(M)$ coming from the Hodge decomposition $H^{3}(M,\mathbb{C})=H^{2,1}(M)\oplus H^{1,2}(M)$. It is a principally polarized abelian variety of dimension 55 [3]. The intermediate Jacobian plays a central role in the geometry of cubic threefolds (for instance in the Clemens–Griffiths irrationality theorem).
14.2. Fermat cubic
Let
$X_{0}:\quad x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+x_{4}^{3}=0$
be the Fermat cubic threefold. We consider the one-parameter deformation
(14.1) $X_{t}:\quad F_{t}:=x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+x_{4}^{3}-3t\,x_{0}x_{1}x_{2}=0,\qquad t\in U\subset\mathbb{C},$
which is smooth for generic $t$ (i.e. away from the discriminant locus). Write $\pi:\mathcal{X}\to U$ for the corresponding family.
Let $\Omega$ be the standard homogeneous 44-form on $\mathbb{C}^{5}$,
$\Omega=\sum_{i=0}^{4}(-1)^{i}x_{i}\,dx_{0}\wedge\cdots\wedge\widehat{dx_{i}}\wedge\cdots\wedge dx_{4}.$
By Griffiths’ description of primitive cohomology of smooth hypersurfaces, the Hodge bundle $E=\pi_{*}\Omega^{2}_{\mathcal{X}/U}$ can be realized via residues [8, 20]:
$H^{2,1}(X_{t})\ \cong\ \Bigl\{\ \operatorname{Res}_{X_{t}}\frac{P\Omega}{F_{t}^{2}}\ :\ P\in R_{1}(F_{t})\ \Bigr\},$
where $R_{\bullet}(F_{t})$ denotes the Jacobian ring of $F_{t}$. In our cubic threefold situation one has $R_{1}(F_{t})\cong\mathbb{C}\langle x_{0},\dots,x_{4}\rangle$, hence a convenient holomorphic frame is given by the residue classes
(14.2) $\omega_{i}(t):=\operatorname{Res}_{X_{t}}\frac{x_{i}\,\Omega}{F_{t}^{2}},\qquad i=0,1,2,3,4.$
We write $\omega=(\omega_{0},\dots,\omega_{4})^{t}$.
14.3. Canonical second-order equation
The Gauss–Manin connection on $H^{3}_{\mathrm{dR}}(\mathcal{X}/U)$ yields a first order system of rank $10$. As in Part IV, under a genericity hypothesis (Definition 12.1) one can pass canonically to a second order equation on $E$ with coefficients in $\operatorname{End}(E)$:
(14.3) $\omega^{\prime\prime}=2A(t)\,\omega^{\prime}+q(t)\,\omega,\qquad A(t),q(t)\in\operatorname{End}(E_{t}),$
where primes denote differentiation with respect to $t$. Concretely, differentiating (14.2) produces residue classes with higher pole order; Griffiths–Dwork reduction expresses these classes back in the $\mathbb{C}$-span of $\{\omega_{0},\dots,\omega_{4},\omega_{0}^{\prime},\dots,\omega_{4}^{\prime}\}$. This reduction procedure is standard in the theory of Picard–Fuchs equations for hypersurfaces; see [4, 7]. We omit the reduction computations and focus on the structural features of (14.3) relevant to our non-abelian Schwarzian viewpoint.
The family (14.1) is invariant under a finite diagonal symmetry group $G$ generated by
$(x_{0},x_{1},x_{2},x_{3},x_{4})\longmapsto(\zeta_{0}x_{0},\zeta_{1}x_{1},\zeta_{2}x_{2},\zeta_{3}x_{3},\zeta_{4}x_{4}),\qquad\zeta_{i}^{3}=1,\ \zeta_{0}\zeta_{1}\zeta_{2}=1.$
This group acts fiberwise on $X_{t}$ and hence on $H^{2,1}(X_{t})$, and it commutes with the Gauss–Manin connection. In particular, the coefficient matrices $A(t)$ and $q(t)$ in (14.3) commute with the $G$-action.
The residue classes (14.2) are eigenvectors for the diagonal action. Indeed, $F_{t}$ is $G$-invariant, and $\Omega$ scales by $\zeta_{0}\zeta_{1}\zeta_{2}\zeta_{3}\zeta_{4}=\zeta_{3}\zeta_{4}$, so $\omega_{i}$ transforms by a one-dimensional character of $G$. Consequently, the $G$-representation on $E_{t}\simeq H^{2,1}(X_{t})$ splits as a direct sum of five one-dimensional character spaces. By Schur’s lemma, any endomorphism commuting with $G$ preserves each character line. Therefore, in a $G$-eigenbasis the matrices $A(t)$ and $q(t)$ are diagonal:
(14.4) $A(t)=\mathrm{diag}(a_{0}(t),\dots,a_{4}(t)),\qquad q(t)=\mathrm{diag}(q_{0}(t),\dots,q_{4}(t)),$
for meromorphic functions $a_{i}(t),q_{i}(t)$ on $U$ (with poles along the discriminant). Equivalently, the system (14.3) decouples into five scalar second order ODEs
$\omega_{i}^{\prime\prime}=2a_{i}(t)\,\omega_{i}^{\prime}+q_{i}(t)\,\omega_{i},\qquad i=0,1,2,3,4.$
For the deformation (14.1), the diagonal coefficients can be made completely explicit.
Proposition 14.1.
For the family (14.1), in the residue frame (14.2) the second-order equation (14.3) has diagonal coefficient matrices
$A(t)=\mathrm{diag}(0,0,0,a(t),a(t)),\qquad q(t)=\mathrm{diag}(0,0,0,b(t),b(t)),$
where
(14.5) $a(t)=\frac{3t^{2}}{2(1-t^{3})},\qquad b(t)=\frac{t}{1-t^{3}}.$
Equivalently,
$\omega_{0}^{\prime}=\omega_{1}^{\prime}=\omega_{2}^{\prime}=0,\qquad\omega_{3}^{\prime\prime}=\frac{3t^{2}}{1-t^{3}}\,\omega_{3}^{\prime}+\frac{t}{1-t^{3}}\,\omega_{3},\qquad\omega_{4}^{\prime\prime}=\frac{3t^{2}}{1-t^{3}}\,\omega_{4}^{\prime}+\frac{t}{1-t^{3}}\,\omega_{4}.$
In the more classical notation $\omega^{\prime\prime}=P(t)\,\omega^{\prime}+Q(t)\,\omega$ we have $P(t)=2A(t)$ and $Q(t)=q(t)$, hence
$P(t)=\mathrm{diag}\Bigl(0,0,0,\frac{3t^{2}}{1-t^{3}},\frac{3t^{2}}{1-t^{3}}\Bigr),\qquad Q(t)=\mathrm{diag}\Bigl(0,0,0,\frac{t}{1-t^{3}},\frac{t}{1-t^{3}}\Bigr).$
Proof.
We sketch the standard reduction argument and omit the routine algebra. Differentiating (14.2) with respect to $t$ gives
$\omega_{i}^{\prime}\;=\;-2\,\operatorname{Res}_{X_{t}}\frac{x_{i}\,(\partial_{t}F_{t})\,\Omega}{F_{t}^{3}}\;=\;6\,\operatorname{Res}_{X_{t}}\frac{x_{i}\,x_{0}x_{1}x_{2}\,\Omega}{F_{t}^{3}}.$
For $i=0,1,2$ the corresponding degree–44 numerator lies in the Jacobian ideal of $F_{t}$, hence its residue class vanishes; this yields $\omega_{0}^{\prime}=\omega_{1}^{\prime}=\omega_{2}^{\prime}=0$. For $i=3,4$ one differentiates once more, obtaining a residue with pole order 44 whose degree–77 numerator is a multiple of $(x_{0}x_{1}x_{2})^{2}$. Using the relations $\partial_{x_{0}}F_{t}=3(x_{0}^{2}-tx_{1}x_{2})$, $\partial_{x_{1}}F_{t}=3(x_{1}^{2}-tx_{0}x_{2})$, $\partial_{x_{2}}F_{t}=3(x_{2}^{2}-tx_{0}x_{1})$, together with the Griffiths–Dwork reduction identity [4, 7], one rewrites this pole–44 class in the span of $\omega_{i}$ and $\omega_{i}^{\prime}$. In the present monomial deformation, the reduction is particularly simple and produces the diagonal coefficients (14.5). ∎
For $e=\tfrac{1}{2}$ the matrix Schwarzian attached to (14.3) is
$\mathcal{S}_{A,q}=2(F_{A}-q),\qquad F_{A}=A^{\prime}-A^{2}.$
In the diagonal frame (14.4), $\mathcal{S}_{A,q}$ is diagonal with entries $2(a_{i}^{\prime}-a_{i}^{2}-q_{i})$; thus the matrix Schwarzian packages scalar Schwarzian expressions into a single $\operatorname{End}(E)$-valued invariant.
For the explicit monomial deformation (14.1), combining Proposition 14.1 with the definition of curvature yields a closed form for the matrix Schwarzian:
(14.6) $\mathcal{S}_{A,q}(t)=\mathrm{diag}(0,0,0,s(t),s(t)),\qquad s(t)=\frac{t(t^{3}+8)}{2(1-t^{3})^{2}}.$
More invariantly, by the discussion in Section 7 and Corollary 8.10, the conjugacy class of $\mathcal{S}_{A,q}$ is independent of the choice of a fundamental solution of the second order system, and the functions
$\operatorname{tr}(\mathcal{S}_{A,q}^{r}),\qquad 1\leq r\leq 5,$
define intrinsic invariants of the one-parameter family (14.1). In the explicit diagonal form (14.6) we have
$\operatorname{tr}(\mathcal{S}_{A,q}^{r})(t)=2\,s(t)^{r}=2\left(\frac{t(t^{3}+8)}{2(1-t^{3})^{2}}\right)^{r},\qquad 1\leq r\leq 5.$
In this way, periods of cubic threefolds give a higher-dimensional test case for the non-abelian Schwarzian formalism, analogous to Dedekind’s genus-one picture but now in weight three and with genuinely non-abelian coefficient objects.
Part V The Mass-Spring System
In this part we discuss mass–spring systems as a mechanical testing ground for the non-abelian Schwarzian, and use them to motivate the terminology “classical” and “quantum” connections.
15. The mass–spring system
The mass–spring system provides a concrete source of second-order equations with matrix coefficients. It also serves as a simple model for the local-to-global viewpoint underlying our quantum formalism: the coefficients are best regarded as geometric objects on a time curve rather than as functions of an absolute time parameter.
15.1. What is time?
Imagine a universe in which time runs on an abstract Riemann surface $X$: a universe without a beginning, without an end, without any distinguished moment, and without any “bangs.” A viewpoint of this kind appears in no–boundary proposal of Hartle–Hawking [13]. In that proposal, the question of what happened before the Big Bang becomes as meaningless as asking what lies south of the South Pole.
A holomorphic time–chart on $X$ may be thought of as a local clock $t$, while another chart provides a clock $\tau$. The two clocks are related by a biholomorphic “time difference” $\lambda$ satisfying $t=\lambda(\tau)$.
15.2. Classical mass–spring systems
We begin with a frictionless unit mass attached to a spring of stiffness $k>0$. In the usual sign convention the displacement $\psi$ satisfies $\ddot{\psi}+k\psi=0$. To match the normalization used throughout the paper, we rewrite this as
(15.1) $\ddot{\psi}=q\,\psi,\qquad q=-k.$
$k_{1}$$k_{2}$$\psi_{1}$$\psi_{2}$
Figure 1. Mass–spring system with $n=2$.
For a system with two degrees of freedom ($n=2$), with spring constants $k_{1}$ and $k_{2}$ , one obtains the coupled system
(15.2) $\ddot{\psi}_{1}$ $=-(k_{1}+k_{2})\,\psi_{1}+k_{2}\,\psi_{2},$
$\ddot{\psi}_{2}$ $=k_{2}\,\psi_{1}-k_{2}\,\psi_{2}.$
Equivalently, writing $\boldsymbol{\psi}=\begin{pmatrix}\psi_{1}\\ \psi_{2}\end{pmatrix}$ this is a $2\times 2$ second-order matrix ODE
(15.3) $\ddot{\boldsymbol{\psi}}=q\,\boldsymbol{\psi},\qquad q=\begin{pmatrix}-k_{1}-k_{2}&k_{2}\\ k_{2}&-k_{2}\end{pmatrix}$.
In the presence of friction (damping) the system takes the form
(15.4) $\ddot{\boldsymbol{\psi}}=p\,\dot{\boldsymbol{\psi}}+q\,\boldsymbol{\psi},$
where $p$ is the damping matrix and $q$ is the stiffness matrix.
Now imagine that $p$ and $q$ vary meromorphically with respect to (complex) time. In a local clock $t$ we then have
(15.5) $\ddot{\boldsymbol{\psi}}=p(t)\,\dot{\boldsymbol{\psi}}+q(t)\,\boldsymbol{\psi},\qquad p(t),q(t)\in M_{n}(\mathbb{C}).$
15.3. Eternal mass–spring system
Call such a system eternal if it is compatible with changes of local clock, i.e. if it depends only on the time curve $X$ and not on the chosen coordinate. Concretely, under a change of clock $t=\lambda(\tau)$, the chain rule above shows that (15.5) retains the same form provided the coefficients transform by
$\widetilde{p}(\tau)=\lambda^{\prime}(\tau)\,\bigl(p\circ\lambda\bigr)(\tau)+\frac{\lambda^{\prime\prime}(\tau)}{\lambda^{\prime}(\tau)}\,I,\qquad\widetilde{q}(\tau)=(\lambda^{\prime}(\tau))^{2}\,\bigl(q\circ\lambda\bigr)(\tau),$
where II denotes the identity matrix. In particular, $p$ transforms affinely (with an inhomogeneous $\lambda^{\prime\prime}/\lambda^{\prime}$ term), while $q$ transforms tensorially with weight 22. In our notation, for an eternal system the quantity $A=p/2$ transforms from one chart to another as a time-connection with eccentricity $e=\tfrac{1}{2}$, while $q$ transforms as a time-quadratic differential.
A natural coordinate-free invariant is obtained from the Schwarzian curvature. For $e=\tfrac{1}{2}$ the curvature of AA is
$F_{A}=A^{\prime}-A^{2},$
and the associated Schwarzian is $S_{A,q}:=F_{A}-q$. While $S_{A,q}$ has a scalar projective anomaly, its traceless part does not. Consequently the projective curvature
(15.6) $\mathcal{R}:=S_{A,q}-\frac{1}{n}\operatorname{tr}(S_{A,q})\,I$
is a well-defined, eternal $\operatorname{End}(\mathbb{C}^{n})$-valued meromorphic quadratic differential on $X$, independent of the local clock. Notice that $\mathcal{R}$ always vanishes when $n=1$.
15.4. Quantum mass–spring systems
In the quantum mass–spring system one replaces the matrix-valued coefficients $p$ and $q$ by local operators acting on the vector space of local holomorphic functions. Under changes of local clock these operators are required to satisfy the quantum transformation laws of Definitions 5.1 and 5.3, so that the damping term defines a quantum time-connection and the stiffness term defines a quantum time-quadratic differential. From this viewpoint, a specific moment plays no distinguished role; what matters is how the local operators glue together globally on the time curve $X$.
This resonates with a point-free theme in Grothendieck’s writings: the structure of a space is encoded by the sheaf-theoretic data that glue on overlaps, and “points” are secondary, recovered only when needed from the function theory [10]. Quantum theory exhibits an analogous shift of emphasis, replacing a classical phase point by an algebra of observables. Our operator-valued coefficients, together with their cocycle laws under changes of clock, fit naturally into this algebraic viewpoint and help explain the terminology “quantum” in this paper.